### v5

# Recovering contact forms from boundary data

Abstract.
Let $X$ be a compact connected smooth manifold with boundary. The paper deals with contact $1$-forms $\beta$ on $X$, whose Reeb vector fields $v_{\beta}$ admit Lyapunov functions $f$.
We tackle the question: how to recover $X$ and $\beta$ from the appropriate data along the boundary $\partial X$? We describe such boundary data and prove that they allow for a reconstruction of the pair $(X,\beta)$, up to a diffeomorphism of $X$. We use the term “holography” for the reconstruction. We say that objects or structures inside $X$ are holographic, if they can be reconstructed from their $v_{\beta}$-flow induced “shadows” on the boundary $\partial X$.
We also introduce numerical invariants that measure how “wrinkled” the boundary $\partial X$ is with respect to the $v_{\beta}$-flow and study their holographic properties under the contact forms preserving embeddings of equidimensional contact manifolds with boundary. We get some “non-squeezing results” about such contact embedding, which are reminiscent of Gromov’s non-squeezing theorem in symplectic geometry.
1. Introduction
In this paper, we study contact $1$-forms $\beta$ and contact structures $\xi_{\beta}$ they generate on compact connected smooth $(2n+1)$-manifolds $X$ with boundary. We focus on residual structures along the boundary $\partial X$ that allow to reconstruct $X$ and the contact form $\beta$ or the contact structure $\xi_{\beta}$ on $X$, up to a smooth diffeomorphism. When possible, such a reconstruction deserves the name of “holography”.
We understand better the reconstruction of contact forms $\beta$ than the reconstruction of associated contact structures $\xi_{\beta}$ from the appropriate boundary data. Thus, our results mainly focus on contact forms. In fact, modulo the Holography Theorem from [K4], our techniques are quite pedestrian, but the questions we tackle here seem to be novel. Some ideas of this paper can be traced back to [K2] and [K5].
The earlier version of this paper [K6] contains a wrong claim (Lemma 6.1). In the present version, we have deleted the old Section 6, Proposition 7.2, and Corollary 7.2, which depended on that erroneous lemma. Also the present Conjecture 4.2 was stated in the earlier versions as Theorem 4.1.
We tried to make this text self-contained and each section readable almost independently from the others.
Let us describe the general plan of the paper and some of its results informally. The paper consists of the following seven sections:
  1. (1)
    Introduction.
  2. (2)
    Traversing, boundary generic, and traversally generic vector fields.
    This section contains needed trivia about vector fields on manifolds with boundary, based on [K1] and [K3].
  3. (3)
    Holography on manifolds with boundary and causality maps.
    This section discusses Holographic properties of general compact manifolds with boundary, equipped with traversing vector fields $v$, and the causality maps $C_{v}$, generated by $v$ (see also [K4], [K2], [K5] for more details).
  4. (4)
    Contact Holography: recovering contact geometry from the boundary data.
    We start by exhibiting differences between generic non-vanishing vector fields and Reeb vector fields, produced by contact forms (see Propositions 4.1, 4.2).
    The main results of this section are the holography Theorem 4.1, together with its Corollary 4.8. They claim that certain list $\{C_{v_{\beta}},\beta^{\partial},(f^{\bullet})^{\partial}\}$ of boundary-confined data is sufficient for a reconstruction of the bulk $X$ and a contact form $\beta$ on it, up to a diffeomorphism $\Phi:X\to X$ which is the identity on the boundary $\partial X$. These results lead to Theorem 4.2 claiming the existence of a group-theoretical section $\sigma:\mathsf{Diff}^{\infty}_{+}(\partial X;\,C_{v_{\beta}},\beta^{\partial},(f^{\bullet})^{\partial})\to\mathsf{Diff}^{\infty}_{+}(X;\beta,f^{\bullet})$ with values in the group of $\beta$-contactomorphisms $\mathsf{Diff}^{\infty}_{+}(X;\beta,f^{\bullet})$ (see Section 4 for the definitions of all these decorations).
  5. (5)
    On contact forms and contact vector fields on manifolds with boundary.
    The main results of this section is Corollary 5.2 dealing with special $1$-parameter families of contact forms in the spirit of Gray Stability Theorem [Gray]. In this section, we analyze the restrictions imposed on contact fields and their flows by the existence of flow-invariant boundary $\partial X$ (see Corollary 5.1).
  6. (6)
    Non-squeezing of $2n$-symplectic volumes, Morse wrinkles of Reeb fields, and holography.
    This section contains the “non-squeezing” Theorem 6.2. It imposes constraints on contact embeddings $\Psi$ of $(2n+1)$-compact manifolds with boundary into a given contact $(2n+1)$-manifold, which carries a traversing Reeb vector field. It is an analogue of Gromov’s non-squeezing Theorem [Gr2] in Symplectic Geometry. The new phenomenon is that the $(2n-1)$-volume of the equatorial locus in the target restricts the $(2n-1)$-volume of the equatorial locus in the source, provided that we restrict a priory the number of times the Reeb trajectories in the target can hit the boundary of the $\Psi$-image.
    Then we introduce numerical invariants of the Reeb flows on contact manifolds with boundary. They are based on volumes of the Morse strata of the boundary, which measure how “wrinkled” the boundary is with respect to the Reeb flow. We study the holographic properties of these invariants under contact embeddings of equidimensional manifolds (Proposition 6.1, Theorem 6.3).
  7. (7)
    Shadows of Legandrian submanifolds of $(X,\beta)$ on the screen $\partial_{1}^{+}X(v_{\beta})$ and on the trajectory space $\mathcal{T}(v_{\beta})$.
    In this section, we consider Legandrian submanifolds $L$ of a given contact compact manifold $X$ whose Reeb vector field $v_{\beta}$ is traversing. Proposition 7.1 describes the structures on the boundary $\partial X$, sufficient for reconstructing $L$. These structures include the $(-v_{\beta})$-flow generated Lagrangian “shadow” $L^{\dagger}\subset\partial X$ of $L$. For sub-Legandrian links, in Corollary 7.2, we give a boundary-localized criterion when the $v_{\beta}$-flow through $L$ interacts with a $v_{\beta}$-concave portion of the boundary $\partial X$.
2. Traversing, boundary generic, & traversally generic vector fields
Let $X$ be a compact connected smooth $(n+1)$-dimensional manifold with boundary.
Definition 2.1.
A vector field $v$ on $X$ is called traversing if each $v$-trajectory is homeomorphic to ether a closed interval, or to a singleton.
We denote by $\mathcal{V}_{\mathsf{trav}}(X)$ the space of all traversing vector fields on $X$. $\diamondsuit$
The $1$-dimensional oriented foliations, generated by such traversing vector fields were originally studied in [EG]. The book [K3] may serve as a general reference source for their properties.
It follows that any traversing vector field does not vanish and is of the gradient type [K1], i.e., there exists a smooth Lyapunov function $f:X\to\mathbb{R}$ such that $df(v)>0$ in $X$. Moreover, the converse is true: any non-vanishing gradient-type vector field $v$ is traversing.
Any smooth vector field $v$ on $X$, which does not vanish along the boundary $\partial X$, gives rise to a partition $\partial_{1}^{+}X(v)\cup\partial_{1}^{-}X(v)$ of the boundary $\partial_{1}X=_{\mathsf{def}}\partial X$ into two sets: the locus $\partial_{1}^{+}X(v)$, where the field is directed inward of $X$ or is tangent to $\partial X$, and the locus $\partial_{1}^{-}X(v)$, where it is directed outward of $X$ or is tangent to $\partial X$.
We assume that $v|_{\partial X}$, viewed as a section of the quotient line bundle $T(X)/T(\partial X)$ over $\partial X$, is transversal to its zero section. This assumption implies that both sets, $\partial^{+}_{1}X(v)$ and $\partial^{-}_{1}X(v)$, are compact manifolds which share a common smooth boundary $\partial_{2}X(v)=_{\mathsf{def}}\partial(\partial_{1}^{+}X(v))=\partial(\partial_{1}^{-}X(v))$. Evidently, $\partial_{2}X(v)$ is the locus where $v$ is tangent to the boundary $\partial X$.
In his groundbreaking work [Mo], M. Morse made an important observation: for a generic vector field $v$, the tangent locus $\partial_{2}X(v)$ inherits a similar structure in connection to $\partial_{1}^{+}X(v)$, as $\partial_{1}X$ has in connection to $X$. That is, $v$ gives rise to a partition $\partial_{2}^{+}X(v)\cup\partial_{2}^{-}X(v)$ of $\partial_{2}X(v)$ into two sets: the locus $\partial_{2}^{+}X(v)$, where the vector field $v$ is directed inward of $\partial_{1}^{+}X(v)$ or is tangent to $\partial_{2}X(v)$, and $\partial_{2}^{-}X(v)$, where it is directed outward of $\partial_{1}^{+}X(v)$ or is tangent to $\partial_{2}X(v)$. Again, we assume that the restriction $v|_{\partial_{2}X(v)}$, viewed as a section of the quotient line bundle $T(\partial X)/T(\partial_{2}X(v))$ over $\partial_{2}X(v)$, is transversal to its zero section.
For generic fields, this structure replicates itself: the cuspidal locus $\partial_{3}X(v)$ is defined as the locus where $v$ is tangent to $\partial_{2}X(v)$; $\partial_{3}X(v)$ is divided into two manifolds, $\partial_{3}^{+}X(v)$ and $\partial_{3}^{-}X(v)$. In $\partial_{3}^{+}X(v)$, the field is directed inward of $\partial_{2}^{+}X(v)$ or is tangent to its boundary, in $\partial_{3}^{-}X(v)$, outward of $\partial_{2}^{+}X(v)$ or is tangent to its boundary. We can repeat this construction until we reach the zero-dimensional stratum $\partial_{n+1}X(v)=\partial_{n+1}^{+}X(v)\cup\partial_{n+1}^{-}X(v)$.
To achieve some uniformity in our notations, put $\partial_{0}^{+}X=_{\mathsf{def}}X$ and $\partial_{1}X=_{\mathsf{def}}\partial X$.
Thus, a generic vector field $v$ on $X$ gives rise to two stratifications:
$\partial X=_{\mathsf{def}}\partial_{1}X\supset\partial_{2}X(v)\supset\dots\supset\partial_{n+1}X(v),$
(2.1) $X=_{\mathsf{def}}\partial_{0}^{+}X\supset\partial_{1}^{+}X(v)\supset\partial_{2}^{+}X(v)\supset\dots\supset\partial_{n+1}^{+}X(v),$
the first one by closed submanifolds, the second one—by compact ones. Here $\dim(\partial_{j}X(v))=\dim(\partial_{j}^{+}X(v))=n+1-j$.
We will often use the notation “$\partial_{j}^{\pm}X$” instead of “$\partial_{j}^{\pm}X(v)$” when the vector field $v$ is fixed or its choice is obvious.
These considerations motivate a more formal definition.
Definition 2.2.
Let $X$ be a compact smooth $(n+1)$-dimensional manifold with boundary $\partial X\neq\emptyset$, and $v$ a smooth vector field on $X$.
We say that $v$ is boundary generic if the restriction $v|_{\partial X}$ does not vanish and produces a filtrations of $X$ as in (2). Its strata $\{\partial_{j}^{+}X\subset\partial_{j}X\}_{1\leq j\leq n+1}$ are defined inductively in $j$ as follows:
  • $\partial_{0}X=_{\mathsf{def}}\partial X$, $\partial_{1}X=_{\mathsf{def}}\partial X$ (hence $\partial_{0}X$ and $\partial_{1}X$—the base of induction—do not depend on $v$).
  • for each $k\in[1,j]$, the $v$-generated stratum $\partial_{k}X$ is a closed smooth submanifold of $\partial_{k-1}X$,
  • the field $v$, viewed as section of the quotient 1-bundle
    $T_{k}^{\nu}=_{\mathsf{def}}T(\partial_{k-1}X)/T(\partial_{k}X)\longrightarrow\partial_{k}X,$
    is transversal to the zero section of $T_{k}^{\nu}\to\partial_{k}X$ for all $k\leq j$.
  • the stratum $\partial_{j+1}X$ is the zero set of the section $v\in T_{j}^{\nu}$.
  • the stratum $\partial^{+}_{j+1}X\subset\partial_{j+1}X$ is the locus where $v$ points inside of $\partial_{j}^{+}X$.
We denote the space of boundary generic vector fields on $X$ by the symbol $\mathcal{B}^{\dagger}(X)$. $\diamondsuit$
For the rest of the paper, we assume that the field $v$ on $X$ extends to a vector field $\hat{v}$ on some open manifold $\hat{X}$ which properly contains $X$. We treat the extension $(\hat{X},\hat{v})$ as a germ $Op(X,v)$ that contains $(X,v)$. One may think of $\hat{X}$ as being obtained from $X$ by attaching an external collar to $X$ along $\partial_{1}X$. In fact, the treatment of $(X,v)$ will not depend on the germ of extension $(\hat{X},\hat{v})$, but many constructions are simplified by introducing an extension.
The trajectories $\gamma$ of a boundary generic vector field $v$ on $X$ interact with the boundary $\partial_{1}X$ so that each point $a\in\gamma\cap\partial_{1}X$ acquires a multiplicity $m(a)\in\mathbb{N}$, the order of tangency of $\gamma$ to $\partial_{1}X$ at $a$. We associate a divisor $D_{\gamma}=\sum_{a\,\in\,\gamma\cap\partial_{1}X}m(a)\cdot a$ with each $v$-trajectory $\gamma$. In fact, for any boundary generic $v$, $m(a)\leq\dim(X)$ and the support of $D_{\gamma}$ is finite ([K2], Lemma 3.1).
Therefore, we may associate also a finite ordered sequence $\omega(\gamma)=(\omega_{1},\omega_{2},\dots,\omega_{q})$ of multiplicities with each $v$-trajectory $\gamma$. The multiplicity $\omega_{i}$ is the order of tangency between the curve $\gamma$ and the hypersurface $\partial_{1}X$ at the $i^{th}$ point of the finite set $\gamma\cap\partial_{1}X$. The linear order in $\gamma\cap\partial_{1}X$ is determined by $v$.
For each trajectory $\gamma$ of a boundary generic and traversing $v$, we introduce two important quantities:
(2.2) $m(\gamma)=_{\mathsf{def}}\sum_{a\in\gamma\cap\partial_{1}X}\;m(a),\;\text{ and }\;m^{\prime}(\gamma)=_{\mathsf{def}}\sum_{a\in\gamma\cap\partial_{1}X}\;(m(a)-1),$
the multiplicity and the reduced multiplicity of the trajectory $\gamma$.
Similarly, for a sequence $\omega=(\omega_{1},\omega_{2},\,\dots\,,\omega_{q})$, we introduce its norm and its reduced norm by the formulas:
(2.3) $|\omega|=_{\mathsf{def}}\sum_{i}\;\omega_{i}\quad\text{and}\quad|\omega|^{\prime}=_{\mathsf{def}}\sum_{i}\;(\omega_{i}-1).$
Note that $q$, the cardinality of the support of $\omega$, is equal to $|\omega|-|\omega|^{\prime}$.
We consider also an important subclass of traversing and boundary generic fields $v$, which we call traversally generic (see Definition 2.3 below and Definition 3.2 from [K2]). Such vector fields admit $v$-flow-adjusted local coordinate systems, in which the boundary is given by a quite special polynomial equations, and all the trajectories are parallel to the preferred coordinate axis (see [K2], Lemma 3.4). More importantly, for such vector fields, the smooth topological type of the flow in the vicinity of each trajectory of a combinatorial type $\omega$ is determined by $\omega$!
Let us give a definition of a traversally generic vector field. Given a boundary generic and traversing vector field $v$, for each trajectory $\gamma$, consider the finite set $\gamma\cap\partial_{1}X=\{a_{i}\}_{i}$ and the collection of tangent spaces $\{T_{a_{i}}(\partial_{j_{i}}X^{\circ})\}_{i}$ to the pure strata $\{\partial_{j_{i}}X^{\circ}\}_{i}$. Note that each point $a_{i}$ belongs to a unique stratum with the maximal possible index $j_{i}:=j(a_{i})$. In fact, each space $T_{a_{i}}(\partial_{j_{i}}X^{\circ})$ is transversal to the curve $\gamma$ [K2].
Let $S$ be a local transversal section of the $\hat{v}$-flow at a point $a_{\star}\in\gamma$, and let $\mathsf{T}_{\star}$ be the space tangent to the section $S$ at $a_{\star}$. Each space $T_{a_{i}}(\partial_{j}X^{\circ})$, with the help of the $\hat{v}$-flow, determines a vector subspace $\mathsf{T}_{i}=\mathsf{T}_{i}(\gamma)$ in $\mathsf{T}_{\star}$. It is the image of the tangent space $T_{a_{i}}(\partial_{j}X^{\circ})$ under the composition of two maps:
(1) the differential of the $v$-flow-generated diffeomorphism that maps $a_{i}$ to $a_{\star}$, and
(2) the linear projection $T_{a_{\star}}(X)\to\mathsf{T}_{\star}$ whose kernel is generated by $v(a_{\star})$.
The configuration $\{\mathsf{T}_{i}\}$ of affine subspaces $\mathsf{T}_{i}\subset\mathsf{T}_{\star}$ is called generic (or stable) when all the multiple intersections of spaces from the configuration have the least possible dimensions, consistent with the dimensions of $\{\mathsf{T}_{i}\}$. In other words,
$\textup{codim}(\bigcap_{s}\mathsf{T}_{i_{s}},\mathsf{T}_{\star})=\sum_{s}\textup{codim}(\mathsf{T}_{i_{s}},\mathsf{T}_{\star})$
for any subcollection $\{\mathsf{T}_{i_{s}}\}$ of spaces from the list $\{\mathsf{T}_{i}\}$.
Consider the case when $\{\mathsf{T}_{i}\}$ are vector subspaces of $\mathsf{T}_{\star}$. If we interpret each subspace $\mathsf{T}_{i}$ as the kernel of a linear epimorphism $\Phi_{i}:\mathsf{T}_{\star}\to\mathbb{R}^{n_{i}}$, then the property of $\{\mathsf{T}_{i}\}$ being generic can be reformulated as the property of the direct product map $\prod_{i}\Phi_{i}:\mathsf{T}_{\star}\to\prod_{i}\mathbb{R}^{n_{i}}$ being an epimorphism. In particular, for a generic configuration of affine subspaces, if a point belongs to several $\mathsf{T}_{i}$’s, then the sum of their codimensions $n_{i}$ does not exceed the dimension of the ambient space $\mathsf{T}_{\star}$.
The definition below resembles and is inspired by the “Condition NC” imposed on, so called, Boardman maps (see [Bo]) between smooth manifolds (see [GG], page 157, for the relevant definitions). In fact, for a generic traversing vector fields $v$, the $v$-flow delivers germs of Boardman maps $p(v,\gamma):\partial_{1}X\to\mathbb{R}^{n}$, available in the vicinity of each trajectory $\gamma$ [K2]. Here $\mathbb{R}^{n}$ is identified with a transversal section $S$ of the flow in the vicinity of $\gamma$.
Definition 2.3.
A traversing vector field $v$ on $X$ is called traversally generic if:
  • the field is boundary generic in the sense of Definition 2.2 ,
  • for each $v$-trajectory $\gamma\subset X$,the collection of subspaces $\{\mathsf{T}_{i}(\gamma)\}_{i}$ is generic in $\mathsf{T}_{\star}$; that is, the obvious quotient map $\mathsf{T}_{\star}\to\prod_{i}\big{(}\mathsf{T}_{\star}/\mathsf{T}_{i}(\gamma)\big{)}$ is surjective.
We denote by $\mathcal{V}^{\ddagger}(X)$ the space of all traversally generic vector fields on $X$. $\diamondsuit$
In fact, $\mathcal{V}^{\ddagger}(X)$ is an open and dense in the space of all traversing vector fields on $X$ [K1].
For a traversally generic vector field $v$ on a $(n+1)$-dimensional $X$, the trajectory space $\mathcal{T}(v)$ is stratified by subspaces, labeled with the elements $\omega$ of a universal poset $\mathbf{\Omega}^{\bullet}_{|\sim|^{\prime}\leq n}\subset\mathbf{\Omega}^{\bullet}$ that is defined by the constraint $|\omega|^{\prime}\leq n$. It depends only on $\dim(X)=n+1$ (see [K3] for the definition and properties of $\mathbf{\Omega}^{\bullet}_{|\sim|^{\prime}\leq n}$).
3. Holography on manifolds with boundary and causality maps
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold with boundary and $v$ a smooth traversing vector field. Recall that $v$ admits a Lyapunov function $f:X\to\mathbb{R}$ so that $df(v)>0$. We assume that $v$ is boundary generic as in Definition 2.2.
Let $\mathcal{F}(v)$ be the $1$-dimensional oriented foliation, generated by the traversing $v$-flow. Such foliations, under the name short foliations were studied in [EG].
We denote by $\gamma_{x}$ the $v$-trajectory through $x\in X$. Since $v$ is traversing and boundary generic, each $\gamma_{x}$ is homeomorphic either a closed segment, or to a singleton.
Definition 3.1.
We say that a boundary generic and traversing vector field $v$ possesses Property $\mathsf{A}$, if each $v$-trajectory $\gamma$ either is transversal to $\partial_{1}X$ at some point of the set $\gamma\cap\partial_{1}X$, or $\gamma\cap\partial_{1}X$ is a singleton $x$ and $\gamma$ is quadratically tangent to $\partial_{1}X$ at $x$. $\diamondsuit$
A traversing vector field $v$ on $X$ induces a structure of a partially-ordered set $(\partial_{1}X,\succ_{v})$ on the boundary $\partial_{1}X$: for $x,y\in\partial X$, we write $y\succ x$ if the two points lie on the same $v$-trajectory $\gamma$ and $y$ is reachable from $x$ by moving in the $v$-direction.
We denote by $\mathcal{T}(v)$ the trajectory space of $v$ and by $\Gamma:X\to\mathcal{T}(v)$ the obvious projection. For a traversing $v$, $\mathcal{T}(v)$ is a compact space in the topology induced by $\Gamma$. Since for a traversing $v$, any trajectory intersects the boundary, we get that $\mathcal{T}(v)$ is a quotient of $\partial_{1}X$ modulo the relation $\succ_{v}$.
A traversing and boundary generic $v$ gives rise to the causality (scattering) map
(3.1) $C_{v}:\partial_{1}^{+}X(v)\to\partial_{1}^{-}X(v)$
that takes any point $x\in\partial_{1}^{+}X(v)$ to the unique consecutive point $y\in\gamma_{x}\cap\partial_{1}^{-}X(v)$, $y\neq x$, which can be reached from $x$ in the $v$-direction. If no such $y$ is available, we put $C_{v}(x)=x$. We stress that typically $C_{v}$ is a discontinuous map.
We notice that, for any smooth positive function $\lambda:X\to\mathbb{R}_{+}$, we have $C_{\lambda\cdot v}=C_{v}$. Thus, the causality map depends only on the conformal class of a traversing vector field $v$.
In the paper, we will discuss two kinds of related holography problems. The first kind amounts to the question:
To what extend given boundary data are sufficient for reconstructing the unknown bulk $X$ and the traversing $v$-flow on it?
This question may be represented symbolically by the following two diagrams.
[…]
(3.2) $(\partial_{1}X,\;\succ_{v},)\;$${{\longrightarrow}}$ $\;(X,\;\mathcal{F}(v)),$
(3.3) $(\partial_{1}X,\;\succ_{v},\;f^{\partial})\;$${{\longrightarrow}}$ $\;(X,\;\mathcal{F}(v),\;f),$
where $\succ_{v}$ denotes the partial order on boundary, defined by the causality map $C_{v}$, and the symbol “${{\longrightarrow}}$” indicates the unknown ingredients of the diagram.
The second kind of problem is:
Given two manifolds, equipped with traversing flows, and a transformation (a smooth diffeomorphism) $\Phi^{\partial}$ of their boundaries, respecting the relevant boundary data, is it possible to extend $\Phi^{\partial}$ to a transformation $\Phi$ of the manifolds that respects the corresponding flows-generated structures in their interior?
This problem is may be represented by the commutative diagrams:
[…]
$(\partial_{1}X_{1},\;\succ_{v_{1}})\;\stackrel{{\scriptstyle\mathsf{inc}}}{{\longrightarrow}}\;(X_{1},\;\mathcal{F}(v_{1}))$
(3.4) $\quad\downarrow\;\Phi^{\partial}\quad\quad\quad\quad\quad\quad\downarrow\;??\;\;\Phi\quad$
$(\partial_{1}X_{2},\;\succ_{v_{2}})\;\stackrel{{\scriptstyle\mathsf{inc}}}{{\longrightarrow}}\;(X_{2},\;\mathcal{F}(v_{2}))$
$(\partial_{1}X_{1},\;\succ_{v_{1}},\;f_{1}^{\partial})\;\stackrel{{\scriptstyle\mathsf{inc}}}{{\longrightarrow}}\;(X_{1},\;\mathcal{F}(v_{1}),f_{1})$
(3.5) $\quad\downarrow\;\Phi^{\partial}\quad\quad\quad\quad\quad\quad\downarrow\;??\;\;\Phi\quad$
$(\partial_{1}X_{2},\;\succ_{v_{2}},\;f_{2}^{\partial})\;\stackrel{{\scriptstyle\mathsf{inc}}}{{\longrightarrow}}\;(X_{2},\;\mathcal{F}(v_{2}),\;f_{2}).$
These two types of problems come in a big variety of flavors, depending on the more or less rich boundary data and on the anticipated quality of the transformations $\Phi$ (homeomorphisms, $\mathsf{PD}$-homeomorphisms, Hölder homeomorphisms with some control of the Hölder exponent, and diffeomorphisms).
4. Contact Holography: recovering contact geometry from the boundary data
Given a vector $v$ and an exterior $k$-form $\alpha$, in what follows, we denote by $v\,\rfloor\,\alpha$ the $(k-1)$-form $\alpha(v\wedge\sim)$.
We start with a lemma that is not related to Contact Geometry.
Lemma 4.1.
[…]
Proof.
Let $\dim X=d$. By definition of a traversing vector field, there exists a Lyapunov function $h:X\to\mathbb{R}$ such that $dh(v)>0$. We embed properly $X$ in a $d$-dimensional open manifold $\hat{X}$ so that $v$ extends to $\hat{v}$ and $h$ to $\hat{h}$, subject to the constraint $d\hat{h}(\hat{v})>0$ in $\hat{X}$. By shrinking $\hat{X}\supset X$ if necessary, we may also extend $\beta$ to $\hat{\beta}$ in $\hat{X}$ so that $\hat{\beta}(\hat{v})=1$ in the vicinity of $X$ in $\hat{X}$.
Using that each $v$-trajectory $\gamma$ extends to a closed segment of the unique $\hat{v}$-trajectory $\hat{\gamma}$, we construct a tubular cover $\{U_{\alpha}\subset\hat{X}\}_{\alpha}$ of $X$ such that there are diffeomorphisms $\phi_{\alpha}:D^{d-1}\times D^{1}\to U_{\alpha}$ that map each segment $a\times D^{1}$, $a\in D^{d-1}$, to a segment of a $\hat{v}$-trajectory. Moreover, $\phi_{\alpha}(D^{d-1}\times\partial D^{1})\subset\hat{X}\setminus X$. Using compactness of $X$, we choose a finite subcover of the cover $\{U_{\alpha}\}_{\alpha}$. Abusing notations, we denote it $\{U_{\alpha}\}_{\alpha}$.
Let $g_{\alpha}:U_{\alpha}\to\mathbb{R}$ be defined by the formula
(4.1) $g_{\alpha}(\phi_{\alpha}(a\times t))=\int_{\phi_{\alpha}([a\times 0),\;(a\times t])}\hat{\beta}\;=_{\mathsf{def}}\int_{\phi_{\alpha}(a\times 0)}^{\phi_{\alpha}(a\times t)}\hat{\beta},$
where $a\in D^{d-1}$. Note that $dg_{\alpha}|_{\gamma}=\hat{\beta}|_{\gamma}$, where $\gamma=\{\phi_{\alpha}(a\times t)\}_{t}$. Since $\hat{\beta}(\hat{v})=1$, we get $dg_{\alpha}(\hat{v})=1$. The finite cover $\{U_{\alpha}\}_{\alpha}$ gives rise to a finite cover $\{V_{\alpha}\}_{\alpha}$ of the compact trajectory space $\mathcal{T}(v)$. A function $\psi:\mathcal{T}(v)\to\mathbb{R}$ is defined to be smooth, if its pull-back $\Gamma^{\ast}(\psi):X\to\mathbb{R}$ is smooth. We use a smooth partition of unity $\{\psi_{\alpha}:V_{\alpha}\to[0,1]\}_{\alpha}$ to introduce a smooth function $f=_{\mathsf{def}}\sum_{\alpha}\psi_{\alpha}\cdot g_{\alpha}.$ By its construction, $\psi_{\alpha}$ has the property $d\psi_{\alpha}(v)=0$. Therefore,
$df(v)=\sum_{\alpha}(\psi_{\alpha}\cdot dg_{\alpha}(v)+g_{\alpha}\cdot d\psi_{\alpha}(v))=\sum_{\alpha}\psi_{\alpha}=1.$
By a similar formula, any closed interval $[a,b]\subset\gamma$, we get
[…]
(4.2) $=\sum_{\alpha}\psi_{\alpha}\cdot\int_{[a,b]}\beta=\int_{[a,b]}\beta.$
Corollary 4.1.
Any traversing vector field $v$ admits a Lyapunov function $g$ such that $dg(v)=1$.
Proof.
If $f:X\to\mathbb{R}$ is a Lyapunov function for $v$, then the $1$-form $\beta=\frac{df}{df(v)}$ satisfies the hypotheses of Lemma 4.1. By the lemma, there exists a Lypunov function $g:X\to\mathbb{R}$ such that $dg(v)=1$. ∎
Let us recall now a few basic definitions from the field of Contact Geometry.
Definition 4.1.
A coorientable contact structure on a $(2n+1)$-dimensional smooth manifold $X$ is a $2n$-dimensional distribution $\xi$ of cooriented hyperplanes in the tangent bundle $TX$, generated by the kernels of a differential $1$-form $\beta$ on $X$ such that $\beta\wedge(d\beta)^{n}>0$ everywhere. Equivalently, the restriction $(d\beta)^{n}|_{\xi}>0$ everywhere.
An coorientable almost contact structure $\xi$ on a $(2n+1)$-dimensional smooth manifold $X$ is a $2n$-dimensional distribution $\eta$ of cooriented hyperplanes in the tangent bundle $TX$, generated by the kernels of a differential $1$-form $\beta$ on $X$, a nonvanishing vector field $v$ such that $\beta(v)=1$ (so, $v$ is transversal to the distribution $\eta$), and a nondegenerate skew-symmetric $2$-form $\omega$ on $\eta$ (equivalently, a complex structure $J:\eta\to\eta$ such that $J^{2}=-Id$).
$\diamondsuit$
If $\lambda:X\to\mathbb{R}_{+}$ is a positive smooth function, then by a direct computation,
$(\lambda\beta)\wedge(d(\lambda\beta))^{n}=\lambda^{n+1}\beta\wedge\Big{(}d\beta+\frac{d\lambda}{\lambda}\wedge\beta\Big{)}^{n}=\lambda^{n+1}[\beta\wedge(d\beta)^{n}]>0.$
As a hyperplane distribution, a cooriented $\xi$ may be identified with the orthogonal complement to a section of the spherical bundle $\mathcal{S}X\to X$, associated with the tangent bundle $TX$ and having the sphere $S^{2n}$ for its fiber. We denote by $\xi^{\partial}$ the restriction of the distribution $\xi$ to the boundary $\partial_{1}X$ (not to be confused with the intersection $\xi\cap\partial_{1}X$). Since the normal (quotient) $1$-bundle $\nu=TX|_{\partial_{1}X}/\,T(\partial_{1}X)$ is trivial, the distribution $\xi^{\partial}$ is a $(2n)$-dimensional orientable subbundle of the bundle $\nu\oplus T(\partial_{1}X)\to\partial_{1}X$.
Definition 4.2.
A vector field $v_{\beta}$ on $X$ is called the Reeb field for a contact $1$-form $\beta$, if $\beta(v_{\beta})=1$ and $v_{\beta}\in\mathsf{ker}(d\beta)$. $\diamondsuit$
The property $\beta\wedge(d\beta)^{n}\neq 0$ implies that the kernel $\mathsf{ker}(d\beta)$ is $1$-dimensional. Thus, any contact form $\beta$ generates a unique Reeb vector field $v_{\beta}$.
Recall the following well-known fact.
Lemma 4.2.
For a contact structure $\xi$, generated by a contact $1$-form $\beta$, there is a $1$-to-$1$ correspondence between the space $C^{\infty}(X)$ of smooth functions and the space $\mathcal{R}eeb(X,\xi)$ of Reeb vector fields amenable to $\xi$. The correspondence is given by the formula $f\longrightarrow v_{e^{f}\cdot\beta}$.
Proof.
Let us fix a contact form $\beta$ such that $\ker(\beta)=\xi$. Then any other contact form, producing $\xi$ can be written as $e^{f}\beta$ for some smooth function $f$. By the definition of Reeb fields, we get $e^{f}\beta(v_{e^{f}\beta})=1=e^{g}\beta(v_{e^{g}\beta})$. If $v_{e^{f}\beta}=v_{e^{g}\beta}$, this is equivalent to the equation $e^{f}=e^{g}$, or to $f=g$. Therefore, the correspondence $f\longrightarrow v_{e^{f}\cdot\beta}$ is a continuous bijection. ∎
The next few lemmas are dealing with the way a contact form $\beta$ and its differential $d\beta$ interact with the boundary $\partial_{1}X$.
Lemma 4.3.
Let $\beta$ be a contact form on a $(2n+1)$-dimensional manifold $X$ and $v_{\beta}$ its boundary generic Reeb vector field.
Then $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}>0$, while $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{-}X(v_{\beta}))}<0$. Thus, $d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}$ is a symplectic $2$-form.
Proof.
Since $\beta\wedge(d\beta)^{n}$ is the volume form on $X$, by the definition of $\partial_{1}^{+}X(v_{\beta})$, we conclude that $v_{\beta}\,\rfloor\,(\beta\wedge(d\beta)^{n})|_{\partial^{+}X(v_{\beta})}>0$ on $\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))$. On the other hand, since $v_{\beta}\,\rfloor\,(d\beta)^{n}=0$,
$v_{\beta}\,\rfloor\,(\beta\wedge(d\beta)^{n})=(v_{\beta}\,\rfloor\,\beta)(d\beta)^{n}-\beta\wedge(v_{\beta}\,\rfloor\,(d\beta)^{n})=(d\beta)^{n}.$
Therefore, $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}>0$. As a result, $d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}$ is a symplectic form. ∎
Corollary 4.2.
Let $\beta$ be a contact form on $(2n+1)$-dimensional $X$, whose Reeb vector field $v_{\beta}$ is boundary generic. Then $\int_{\partial_{2}X(v_{\beta})}\beta\wedge(d\beta)^{n-1}>0$.
Proof.
Since $(d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))})^{n}>0$, by Stokes’ Theorem,
$\int_{\partial_{2}X(v_{\beta})}\beta\wedge(d\beta)^{n-1}=\int_{\partial_{1}^{+}X(v_{\beta})}(d\beta)^{n}>0.$
Lemma 4.4.
Let $v$ be a traversing vector field on an $(2n+1)$-dimensional compact connected oriented smooth manifold $X$ with boundary. Then $X$ admits a $1$-form $\beta$ such that $\beta(v)=1$ and $v\,\rfloor\,d\beta=0$.
Proof.
By definition of a traversing vector field, there exists a Lyapunov function $h:X\to\mathbb{R}$ such that $dh(v)>0$.
As before, we embed properly $X$ in a $(2n+1)$-dimensional open manifold $\hat{X}$ so that $v$ extends to $\hat{v}$ and $h$ to $\hat{h}$, subject to the constraint $d\hat{h}(\hat{v})>0$ in $\hat{X}$. By Corollary 4.1, we may assume that there is another Lyapunov function $\hat{f}:\hat{X}\to\mathbb{R}$ such that $d\hat{f}(\hat{v})=1$ in the vicinity of $X$ in $\hat{X}$.
Using that each $v$-trajectory $\gamma$ extends to a closed segment of the unique $\hat{v}$-trajectory $\hat{\gamma}$, we construct a finite tubular cover $\{U_{\alpha}\subset\hat{X}\}_{\alpha}$ of $X$ as in the proof of Lemma 4.1. The finite cover $\{U_{\alpha}\}_{\alpha}$ gives rise to a finite cover $\{V_{\alpha}\}_{\alpha}$ of the compact trajectory space $\mathcal{T}(v)$.
In the coordinates $\{x_{1},x_{2},\ldots x_{2n}\}$ on $D^{2n}\subset\mathbb{R}^{2n}$, we consider the closed $1$-form
$\rho=_{\mathsf{def}}\;\frac{1}{2}\big{(}\sum_{i=1}^{n}(x_{i+1}dx_{i}-x_{i}dx_{i+1})\big{)}.$
Using the diffeomorphism $\phi_{\alpha}:D^{2n}\times D^{1}\to U_{\alpha}$ as in the proof of Lemma 4.1, we transfer $\rho$ to the $1$-form $\rho_{\alpha}=_{\mathsf{def}}\phi_{\alpha}^{\dagger}(\rho)$ and the $1$-form $dx_{2n+1}$ to $\phi_{\alpha}^{\dagger}(dx_{2n+1})=df$ on $U_{\alpha}$. Put $\beta_{\alpha}=_{\mathsf{def}}df+\rho_{\alpha}$.
Let $\Gamma:X\to\mathcal{T}(v)$ be the obvious projection. Recall that a function $\psi:\mathcal{T}(v)\to\mathbb{R}$ is defined to be smooth, if its pull-back $\Gamma^{\ast}(\psi):X\to\mathbb{R}$ is smooth.
We use a smooth partition of unity $\{\psi_{\alpha}:V_{\alpha}\to[0,1]\}_{\alpha}$ by smooth functions on $\mathcal{T}(v)$ (i.e., by $v$-invariant smooth functions $\psi_{\alpha}$ on $X$) to introduce a smooth $1$-form $\beta=_{\mathsf{def}}\sum_{\alpha}\psi_{\alpha}\beta_{\alpha}$. By its construction, $\psi_{\alpha}$ has the property $d\psi_{\alpha}(v\wedge\sim)=0$. Therefore,
$\beta(v)=\sum_{\alpha}\psi_{\alpha}\cdot\beta_{\alpha}(v)=df(v)\cdot\sum_{\alpha}\psi_{\alpha}=df(v)=1.$
Thus, $\beta$ is a non-vanishing $1$-form. At the same time, since $d(df)=0$, $\psi_{\alpha}(v)=0$, $v\,\rfloor\,d\psi_{\alpha}=0$, and $\rho_{\alpha}(v)=0$, $v\,\rfloor\,d\rho_{\alpha}=0$, we get:
$v\,\rfloor\,d\beta=v\,\rfloor\,d\big{(}\sum_{\alpha}\psi_{\alpha}\cdot\beta_{\alpha}\big{)}=v\,\rfloor\,d\big{(}\sum_{\alpha}\psi_{\alpha}\cdot\rho_{\alpha}\big{)}=$
$=v\,\rfloor\,\big{(}\sum_{\alpha}d\psi_{\alpha}\wedge\rho_{\alpha}\big{)}+v\,\rfloor\,\big{(}\sum_{\alpha}\psi_{\alpha}\,d\rho_{\alpha}\big{)}=0.$
If we would know that $\beta\wedge(d\beta)^{n}>0$ in $X$, we would conclude that $\beta$ is a contact form and $v$ is its Reeb vector field. However, the property $\beta\wedge(d\beta)^{n}>0$ does not hold automatically. Proposition 4.1 spells out the hypotheses that insure the desired property.
Proposition 4.1.
Let $v$ be a traversing boundary generic vector field on a connected compact smooth $(2n+1)$-manifold with boundary. Let $\xi(v)$ be a $(2n)$-subbundle of the tangent bundle $TX$, transversal to $v$. Consider any $1$-form $\beta$ such that $\beta(v)=1$ and $d\beta(v)=0$. (By Lemma 4.4, such a form $\beta$ exists.)
If $\partial_{2}^{-}X(v)=\emptyset$ (i.e., if $v$ is concave), then
  • $\beta$ is a contact form,
  • $v$ is its Reeb vector field $v_{\beta}$,
  • $\xi(v)$ admits a complex/symplectic structure amenable to $d\beta$,
if and only if,
$(d\beta)^{n}\big{|}_{\mathsf{int}(\partial_{1}^{+}X(v))}>0.$
Proof.
We notice that $v\,\rfloor\,(d\beta)=0$ and $v\,\rfloor\,\beta=1$ imply that $\mathcal{L}_{v}\beta=0$. Thus, $\beta$ is $v$-invariant.
By the same two properties of $\beta$, we get
$v\,\rfloor\,(\beta\wedge(d\beta)^{n})=(v\,\rfloor\,\beta)(d\beta)^{n}-\beta\wedge(v\,\rfloor\,(d\beta)^{n})=(d\beta)^{n}.$
Hence, if $(d\beta)^{n}|_{\mathsf{int}(\partial_{1}^{+}X(v))}>0$, then $\big{(}v\,\rfloor\,(\beta\wedge(d\beta)^{n})\big{)}|_{\mathsf{int}(\partial_{1}^{+}X(v))}>0$.
Let $w_{1},w_{2},\ldots w_{2n}$ be a frame in the tangent space $T_{x}\partial_{1}^{+}X(v)$ whose orientation agrees with the orientation of $\partial_{1}X(v)$, induced by the preferred orientation of $X$. Then $v\wedge w_{1}\wedge w_{2}\wedge\ldots\wedge w_{2n}$ is a positive volume. Thus, $(\beta\wedge(d\beta)^{n})(v\wedge w_{1}\wedge w_{2}\wedge\ldots\wedge w_{2n})>0$ on $\mathsf{int}(\partial_{1}^{+}X(v))$. Using that $\beta$ is $v$-invariant, we conclude that $\beta\wedge(d\beta)^{n}$ is a positive volume form along each $v$-trajectory $\gamma$ such that $\gamma\cap\mathsf{int}(\partial_{1}^{+}X(v))\neq\emptyset$. If $\partial_{2}^{-}X(v)=\emptyset$ (this implies that $\partial_{j}X(v))=\emptyset$ for all $j\geq 3$), any trajectory $\gamma$ has this property. Therefore, in such a case, we conclude that $\beta\wedge(d\beta)^{n}>0$ everywhere in $X$. As a result, $\beta$ is a contact form with the Reeb vector field $v_{\beta}=v$ and the bundle $\xi_{\beta}\approx\xi(v)$ admits a symplectic structure, induced by $d\beta$ (equivalently, a complex structure $J:\xi_{\beta}\to\xi_{\beta}$, consistent with $d\beta$).
Conversely, if $\beta$, subject to the properties $\beta(v)=1$ and $d\beta(v)=0$, is such that $\beta\wedge(d\beta)^{n}>0$ in $X$, then $v$ is the Reeb field for $\beta$, and, by Lemma 4.3, $(d\beta)^{n}\big{|}_{\mathsf{int}(\partial_{1}^{+}X(v))}>0$. ∎
Remark 4.1. Proposition 4.1 suggests that should be a geometric connection between cohomology classes with the values in the groups $H^{\ast+1}(X;\pi_{\ast}(SO(2n)/U(n))$ that obstruct the existence of complex (symplectic) structure on $\xi(v)$ and the geometry of the loci in $\partial_{1}^{+}X(v)$, where $(d\beta)^{n}|_{\partial_{1}^{+}X(v)}$ changes sign. Note that this change in sign can be expressed as the global property $\int_{{\partial_{1}^{+}X(v)}}|(d\beta)^{n}|>\int_{{\partial_{1}^{+}X(v)}}(d\beta)^{n}$. $\diamondsuit$
As Corollary 4.5 testifies, not any traversing vector field $v$ is the Reeb field for some contact form $\beta$! Some obstructions for finding such contact forms $\beta$ are of the homotopy-theoretical nature, others (see Proposition 4.2) are more obscure.
At the first glance, it seems probable that any traversing vector field $v$ on a $(2n+1)$dimensional manifold $X$ is a Reeb vector field for a contact form $\beta$, provided that a transversal to $v$ $(2n)$-distribution $\xi$ admits a complex/symplectic structure. However, this conjecture is false, as the next simple example testifies.
[…]
Let us generalize a bit Example 4.1.
Definition 4.3.
A $2k$-shell is the product $S\times[0,1]$, where $S$ is a connected oriented closed smooth manifold of dimension $2k$, $k\leq n$.
Let $v$ be a smooth non-vanishing vector field on a $(2n+1)$-dimensional compact manifold $X$ and let $\xi$ be a $(2n)$-distribution, transversal to $v$ and admitting a complex/symplectic structure.
We say that $(X,v,\xi)$ contains a contact $(2k+1)$-shell if there is a regular embedding $\Psi:S\times[0,1]\to X$ that maps each oriented segment $s\times[0,1]$, $s\in S$, to an oriented segment of a $v$-trajectory and such that a normal $(2n-2k)$-bundle $\nu(\Psi)$ of $\Psi(S\times[0,1])$ in $X$ is a complex/symplectic subbundle of the $2n$-bundle $\xi|_{\Psi(S\times[0,1])}$. $\diamondsuit$
Equivalently, $(X,v,\xi)$ contains a contact $(2k+1)$-shell $\Psi:S\times[0,1]\to X$ if and only if there is a regular embedding $\psi:S\to X$ such that there exists a symplectic $(2n-2k)$-subbundle $\mu$ of $\xi|_{\psi(S)}$ whose fibers are transversal to $\psi(S)$; i.e., $\mathbb{R}_{v}\oplus\mu$ is a normal bundle of $\psi(S)$ in $X$, where $\mathbb{R}_{v}$ stands for the line bundle spanned by $v$.
Proposition 4.2.
If $v_{\beta}$ is the Reeb vector field of a contact form $\beta$ on a compact $(2n+1)$-dimensional manifold $X$, then $(X,v_{\beta},\ker{\beta})$ does not contain any contact $(2k+1)$-shells for all $k\in[1,n]$.
In particular, the sets $\partial_{1}^{\pm}X(v_{\beta})$ have no closed connected components.
Proof.
Let $\beta$ be a contact form whose Reeb field is $v_{\beta}$. Let $\xi_{\beta}=\ker(\beta)$. Assume that $(X,v_{\beta},\xi_{\beta})$ contains a contact $(2k)$-shell $\Psi(S\times[0,1])$. Thanks to Definition 4.3, the normal bundle $\nu(\Psi)$ is a complex/symplectic subbundle of $\xi_{\beta}|_{\Psi(S\times[0,1])}$.
Consider the form $\Theta^{\bullet}=\Psi^{\ast}(\beta\wedge(d\beta)^{k})$ on $S\times[0,1]$. Since $\beta\wedge(d\beta)^{n}$ is a volume form and $(d\beta)^{n-k}|_{\nu(\Psi)}$ is the volume form by Definition 4.3, the form $\Theta^{\bullet}$ must be a volume form on $S\times[0,1]$. Thus, $\beta^{\bullet}=_{\mathsf{def}}\Psi^{\ast}(\beta)$ is a contact form on $S\times[0,1]$ and the transfer $v^{\dagger}$ of $v_{\beta}$ under $\Psi$ is the Reeb vector field for $\beta^{\bullet}$. Moreover, $\partial_{1}^{+}(S\times[0,1])(v^{\dagger})=S\times\{0\}$.
On the other hand, Proposition 4.1 tells us that such $v^{\dagger}$ cannot be the Reeb vector field of the contact form $\beta^{\bullet}$. Indeed, the form $(d\beta^{\bullet})^{k}$, being exact, cannot be positive on the closed manifold $\partial^{+}_{1}(S\times[0,1])(v^{\dagger})=S\times\{0\}$ since $\int_{S}(d\beta^{\bullet})^{k}=\int_{S}\,d\big{(}\beta^{\bullet}\wedge(d\beta^{\bullet})^{k-1}\big{)}=0$. This contradiction proves the proposition. ∎
Corollary 4.3.
Let $\beta$ be a contact form on an oriented smooth $(2n+1)$-manifold $X$. Let the Reeb vector field $v_{\beta}$ of $\beta$ be transversal to a smooth connected oriented hypersurface $Y\subset X$, an open $(2n)$-manifold.
Then $Y$ does not contain any closed smooth oriented $(2k)$-dimensional submanifold $S$, $k\in[1,n]$, whose normal bundle $\nu(S)$ in $Y$ admits a complex/symplectic structure, amenable to the symplectic form $d\beta|_{Y}$.
In particular, this claim is valid for $Y=\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))$.
Proof.
Let $U$ be a smooth regular neighborhood of $Y$ in $X$. Consider the $(-v_{\beta})$ induced projection $Q^{\beta}:U\to Y$. We denote its differential by $Q^{\beta}_{\ast}$. Since $v_{\beta}$ is transversal to $Y$, the map $Q^{\beta}_{\ast}|:\xi_{\beta}|_{Y}\to T(Y)$ is a symplectomorphism with respect to forms $d\beta|_{\xi_{\beta}}$ and $d\beta|_{Y}$. If the normal bundle $\nu(S)$ of $S$ in $Y$ is a symplectic subbundle of $TY$, then $(Q^{\beta}_{\ast})^{-1}\big{(}\nu(S)\big{)}\cap\xi_{\beta}$ is a symplectic subbundle of $\xi_{\beta}|_{(Q^{\beta})^{-1}(S)}$. It is on the level of definitions to check that $(Q^{\beta})^{-1}(S)$ is a contact shell in $U\subset X$. Now the corollary follows from Proposition 4.2. ∎
Motivated by these results and feeling a bit adventurous, let us state the following conjecture.
Conjecture 4.1.
Let $v$ be a non-vanishing (traversing ?) vector on a compact $(2n+1)$-dimensional manifold $X$, such that a normal to $v$ $(2n)$-dimensional distribution $\xi$ admits a complex/symplectic structure. Then $v$ is a Reeb vector field of a contact form on $X$ if and only if $(X,v,\xi)$ does not admit any contact $(2k+1)$-shells for all $k\in[1,n]$. $\diamondsuit$
We denote by $\mathcal{V}(X)$ the space of all vector fields on $X$ in the $C^{\infty}$-topology.
Proposition 4.3.
Let $\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$ be the space of smooth non-vanishing vector fields $v$ on $X$ that admit a transversal to them symplectic $(2n)$-distribution $\xi\subset TX$, where $\dim(X)=2n+1$. Let $k\in[1,n]$.
Then the subset $\mathcal{W}_{k}(X)$ of $\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$, consisting of pairs $(v,\xi)$ that admit contact $(2k+1)$-shells, is open in the $C^{\infty}$-topology.
Therefore, if $\mathcal{W}_{k}(X)\neq\emptyset$, then $\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$ contains an open nonempty set of pairs $(v,\xi)$, where $v$ is not the Reeb vector field for any contact form $\beta$ on $X$, subject to to $(d\beta)^{n}|_{\xi}>0$.
Proof.
Let $(v_{\star},\xi_{\star})\in\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$. Let $\Psi_{\star}:S\times[0,1]\to X$ be a contact $(2k+1)$-shell for the triple $(X,v_{\star},\xi_{\star})$. We denote by $\tilde{\nu}(\Psi_{\star})$ the symplectic subbundle of $\xi_{\star}$ that extends the normal to $\Psi_{\star}(S\times[0,1])$ bundle $\nu(\Psi_{\star})$ into a neighborhood $U_{\star}\subset X$ of $\Psi_{\star}(S\times[0,1])$. Then the set of $v\in\mathcal{V}(X)$, transversal to $\xi_{\star}$, is open. With the manifold $\Psi_{\star}(S\times\{0\})$ and the neighborhood $U_{\star}$ being fixed, each $v$ sufficiently close to $v_{\star}$ defines a new shell $\Psi(v):S\times[0,1]\to X$ by integrating the field $v$, subject to the initial conditions in $\Psi_{\star}(S\times\{0\})$. Indeed, for any $v$ sufficiently close to $v_{\star}$, the symplectic distribution $\tilde{\nu}(\Psi_{\star})$ is still transversal to $\Psi(v)(S\times[0,1])$. ∎
Let $Z$ be a compact smooth $(2n+1)$-dimensional manifold and $f:Z\to\mathbb{R}$ a Morse function. Let $\tilde{v}$ be a $f$-gradient-like vector field on $Z$. We denote by $X$ the complement in $Z$ to a disjoint union of small open $(2n+1)$-balls $\{B_{z}\}$ that are centered on the critical points $\{z\}$ of $f$ and are convex in the local euclidean metric $g_{z}$, adjusted to the local Morse coordinates at $z$. A similar construction makes sense for any closed $1$-form $\alpha$ on $Z$ with Morse type singularities. In their vicinity, $\alpha=df$, where $f$ has the standard Morse form in the Morse coordinates at each singularity. Recall that a vector field $v$ is of the gradient type for $\alpha$ on $X=Z\setminus\bigcup_{z\in\mathsf{sing}(\alpha)}B_{z}$ if $\alpha(v)>0$ on $X$.
Corollary 4.4.
If $Z$ is a closed manifold, then the gradient-like vector field $v$ on the manifold $X=Z\setminus\bigcup_{z\in\mathsf{crit}(f)}B_{z}$ is not a Reeb field of any contact form.
Proof.
Since any Morse function $f:Z\to\mathbb{R}$ on a closed $Z$ has maxima and minima, the spheres $S_{z}=\partial B_{z}$ surrounding such critical points give rise to contact $(2n+1)$-shells. By Proposition 4.2, $v$ cannot be a Reeb vector field. ∎
Corollary 4.5.
Let $Z$ be a compact smooth manifold of an odd dimension that exceeds $7$. Consider a Morse function $f:Z\to\mathbb{R}$ such that $\mathsf{crit}(f)\neq\emptyset$ and its gradient like vector field $v$.
Then $v$ is not a Reeb vector field of any contact form on $X=Z\setminus\bigcup_{z\in\mathsf{crit}(f)}B_{z}.$
Moreover, for any $1$-form $\beta$ on $X$ such that $\beta(v)=1$ and $v\,\rfloor\,d\beta=0$, the form $(d\beta)^{n}\big{|}_{\mathsf{int}(\partial_{1}^{+}X(v))}$ must flip sign.
Proof.
For any $f$-critical point $z$, consider the sphere $S^{2n}_{z}=\partial B_{z}$. Let $a\in H^{2n}(S^{2n}_{z};\mathbb{Z})$ be a generator. Let $\tilde{K}(S^{2n})$ denote the reduced complex $K$-theory of the sphere.
For each complex vector bundle $\eta$ over $S^{2n}$, the $n$-th Chern class $c_{n}(\eta)$ is a multiple of $(n-1)!\,a$. Moreover, for each $m$ such that $m\equiv 0\mod(n-1)!$, there is a unique $[\eta]\in\tilde{K}(S^{2n})$ such that $c_{n}(\eta)=ma$ (see [Huse], 9.8 Corollary).
[…]
$(-1)^{\mathsf{ind}_{z}(f)}\;2\equiv 0\mod(n-1)!\text{ for each }z\in\mathsf{crit}(f).$
Unfortunately, this arithmetic condition can be satisfied only for $n\leq 3$. Therefore, no gradient field $v$ of a Morse function $f:Z\to\mathbb{R}$ can produce a contact form $\beta$ on $X=Z\setminus(\bigcup_{z\in\mathsf{crit}(f)}\;D_{z})$ such that $v_{\beta}=v|_{X}$ in all dimensions $\dim X\geq 9$.
If we identify a small vector $u\in T_{z}Z$ with the point $exp_{g(z)}(u)\in B_{z}$ in the Euclidean metric $g(z)$ in the vicinity of $z$, in terms of the Hessian of $f$ at $z$, we get
$\partial_{1}^{+}X(v)=\coprod_{z\in\mathsf{crit}(f),\;u\in T_{z}Z}\{Hess(f)_{z}(u,u)\geq 0\}\cap\partial B_{z},$
and
$\partial_{2}X(v)=\partial_{2}^{+}X(v)=\coprod_{z\in\mathsf{crit}(f),\;u\in T_{z}Z}\{Hess(f)_{z}(u,u)=0\}\cap\partial B_{z}.$
Therefore, $v$ is concave in relation to $\partial_{1}X$, and Proposition 4.1 is applicable. Hence, by the argument above, for any $1$-form $\beta$ such that $\beta(v)=1$ and $d\beta(v)=0$, the form $(d\beta)^{n}\big{|}_{\mathsf{int}(\partial_{1}^{+}X(v))}$ must flip sign, provided $\dim X\geq 9$. ∎
Corollary 4.5 admits the following strait-forward generalization.
Corollary 4.6.
Let $\tilde{v}$ be a vector field on an odd-dimensional manifold $Y$ such that the zeros of $\tilde{v}$ form a finite set $\mathcal{Z}_{\tilde{v}}$. Let $U(\mathcal{Z}_{\tilde{v}})$ be an open regular neighborhood of $\mathcal{Z}_{\tilde{v}}$ in $Y$. Put $X=Z\setminus U(\mathcal{Z}_{\tilde{v}})$ and let $v=\tilde{v}|_{X}$.
Then a necessary condition for $v$ being a Reeb vector field of a contact form $\beta$ on $X$ is
[…]
where $\mathsf{ind}_{z}(v)$ denotes the local index of the vector field $v$ at $z$. $\diamondsuit$
Conjecture 4.2.
Let $X$ be a compact connected oriented smooth manifold with boundary, $\dim(X)=2n+1$. Let $\beta$ be a contact form on $X$ such that its Reeb vector field $v_{\beta}$ is transversal to some $\hat{v}_{\beta}$-invariant codimension one foliation $\mathcal{G}$ on an open $(2n+1)$-manifold $\hat{X}$ properly containing $X$. Here the vector field $\hat{v}_{\beta}$ on $\hat{X}$ is such that $\hat{v}_{\beta}|_{X}=v_{\beta}$.
Then there is a smooth isotopy of $\{\phi_{t}:X\to X\}_{t\in[0,1]}$ such that the Reeb vector field $v_{\beta_{1}}$ of the contact form $\beta_{1}=\phi_{1}^{\ast}(\beta)$ is traversing on $\phi_{1}(X)\subset X$. $\diamondsuit$
Since a connected manifold with boundary admits a handle decomposition without handles of the top dimension $12n+$ 1, there is no homotopy-theoretical obstruction for extending any almost contact structure (see Definition 4.1) from the base of each handle to its interior (so that the preconditions for applying the $h$-principle are valid). Therefore, by Gromov’s Theorem [Gr1] (see also its enforcement, the fundamental Theorem 1.1 of Borman, Eliashberg, and Murphy [BEM]), the cooriented contact structure $\xi$ on $X$ exists. However, we do not know whether the appropriate transversal to $v_{\beta}$ and $\hat{v}_{\beta}$-invariant foliation $\mathcal{G}$ exists.
For a traversing vector field $v$ on $X$, every $v$-trajectory $\gamma\subset X$ has a neighborhood $\hat{U}(\gamma)\subset\hat{X}$ and a $t$-parametric family of diffeomorphisms $\{\phi^{t}:\hat{U}(\gamma)\to\hat{U}(\gamma)\}_{t\in[-\epsilon,\epsilon]}$, where $\epsilon>0$ depends on $\gamma$, such that $\phi^{0}=\mathsf{id}$, and $\frac{d}{dt}\phi^{t}(a)\big{|}_{t=0}=v(a)$ for all $a\in\hat{U}(\gamma)\cap X$.
Definition 4.4.
Let $v$ be a vector field on $X$. We say that a contact form $\beta$ is $v$-invariant if $(\phi^{t})^{\ast}(\beta)=\beta$ in $\hat{U}(\gamma)\cap X\cap(\phi^{t})(X)$ for all sufficiently small $t$. Equivalently, $\mathcal{L}_{v}\beta=0$.
Similarly, the contact structure $\xi$ is $v$-invariant, if $(\phi^{t})^{\ast}(\beta)=\lambda_{t}\cdot\beta$, where $\lambda_{t}:X\to\mathbb{R}_{+}$ is a smooth positive function (i.e., $(\phi^{t})^{\ast}(\xi)=\xi$ as distributions). $\diamondsuit$
Definition 4.5.
[…]
  • We call contact forms $\beta_{1}$ and $\beta_{2}$ Reeb-conformally-equivalent ($\mathcal{RC}$-equivalent for short), if their Reeb vector fields $v_{1}$ and $v_{2}$ are proportional: i.e., $v_{2}=\lambda\cdot v_{1}$, where $\lambda:X\to\mathbb{R}_{+}$ is a smooth positive function. Similarly, we call contact structures $\xi_{1}$ and $\xi_{2}$ $\mathcal{RC}$-equivalent if they are generated by $\mathcal{RC}$-equivalent contact forms. $\diamondsuit$
Remark 4.1.
Given a contact form $\beta$, if a smooth positive function $\lambda:X\to\mathbb{R}_{+}$ is such that $\mathcal{L}_{v_{\beta}}\lambda=0$, then the form $\beta_{1}=_{\mathsf{def}}\lambda^{-1}\beta$ is $\mathcal{RC}$-equivalent to the form $\beta$. Moreover, the Reeb vector field $v_{\beta_{1}}=\lambda v_{\beta}$. As a result, positive functions from the algebra $C^{\infty}(\mathcal{T}(v_{\beta}))$ of smooth functions on the trajectory space $\mathcal{T}(v_{\beta})$ act on $\beta$ by multiplication and transform it into contact forms $\mathcal{RC}$-equivalent to $\beta$. $\diamondsuit$
Remark 4.2.
Evidently, $\mathcal{RC}$-equivalent contact forms share the dynamics of their Reeb flows. $\diamondsuit$
Lemma 4.5.
Let $\eta:X\to\mathbb{R}$ be a smooth function such that $\mathcal{L}_{v_{\beta}}\eta=d\eta(v_{\beta})=0$. Consider the $1$-form $\beta_{1}=\beta+d\eta$. Then $\beta_{1}$ is $\mathcal{R}$-equivalent to $\beta$.
Proof.
By definition, $d\eta(v_{\beta})=0,v_{\beta}\,\rfloor\,d\beta=0$, and for dimensional reasons $d\eta\wedge(d\beta)^{n}|_{\ker\beta}=0$. Thus, $d\eta\wedge(d\beta)^{n}=0$. Therefore, $\beta_{1}\wedge(d\beta_{1})^{n}=(\beta+d\eta)\wedge(d\beta)^{n}=\beta\wedge(d\beta)^{n}>0$. Also, $\beta_{1}(v_{\beta})=\beta(v_{\beta})+d\eta(v_{\beta})=1$ and, since $d\beta_{1}=d\beta$, we get $v_{\beta}\,\rfloor\,d\beta_{1}=0$. Hence, $v_{\beta_{1}}=v_{\beta}$. ∎
Lemma 4.6.
For a contact form $\beta$ on $X$, any smooth diffeomorphism $\Phi:X\to X$, that preserves the $\mathcal{RC}$-equivalence class of the contact form $\beta$, preserves the Morse stratifications $\{\partial_{j}X(v_{\beta})\}_{j},\{\partial_{j}^{\pm}X(v_{\beta})\}_{j}$ (see (2)) of the boundary.
Proof.
Recall that $\beta$ uniquely defines the Reeb field $v_{\beta}$, and the conformal class of $v_{\beta}$ determines the $v_{\beta}$-generated Morse stratification of the boundary.
If $\Phi^{\ast}(\beta)=\beta_{1}$ and $v_{\beta_{1}}=\lambda\cdot v_{\beta}$ for some smooth function $\lambda:X\to\mathbb{R}_{>0}$ (i.e., $\beta_{1},\beta$ are $\mathcal{RC}$-equivalent), then the two Morse stratifications coincide, since they are defined in terms of applying iteratively the Lie derivatives $\mathcal{L}_{v_{\beta}}$ and $\mathcal{L}_{v_{\beta_{1}}}=\lambda\cdot\mathcal{L}_{v_{\beta}}$ to an auxiliary function $h:\hat{X}\to\mathbb{R}$ that determines the boundary $\partial_{1}X$ [K1]. By definition, such a function $h$ has the three properties:
[…]
We adopt the notation $Op(Y)$ for the vicinity of a topological subspace $Y$ in the ambient space $X$.
Lemma 4.7 states a relative version of the Gray’s theorem [Gray]. Its claim may be found in [EM], page 95. Lemma 5.5 will generalize a bit Lemma 4.7 (whose proof we skip).
Lemma 4.7.
Let $\{\xi_{t}\}_{t\in[0,1]}$ be a family of contact structures on a compact connected manifold $X$ with boundary such that $\xi_{t}=\xi_{0}$ in the vicinity $Op(\partial_{1}X)$ of $\partial_{1}X$ in $X$.
Then there is an isotopy $\phi_{t}:X\to X$, fixed in $Op(\partial_{1}X)$, such that $(\phi^{t})^{\ast}(\xi_{0})=\xi_{t}$ for all $t\in[0,1]$. $\diamondsuit$
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold $X$ with boundary.
We denote by $\mathsf{Cont.forms}(X)$ the space of smooth contact forms $\beta$ on $X$, and by $\mathsf{Cont.forms}(X,\partial_{1}X;Op(\beta^{\partial}))$ the space of smooth contact forms $\beta$ on $X$, such that the section $\beta^{\partial}=\beta|:Op(\partial_{1}X)\to T^{\ast}(X)|_{Op(\partial_{1}X)}$ is fixed in the vicinity $Op(\partial_{1}X)$ of $\partial_{1}X$. Similarly, we denote by $\mathsf{Cont}(X)$ the space of smooth contact structures $\xi$ on $X$, and by $\mathsf{Cont}(X,\partial_{1}X;Op(\xi^{\partial}))$ the space of smooth contact structures $\xi$ on $X$, such that the distribution $\xi\subset T_{\ast}(X)|_{Op(\partial_{1}X)}$ is fixed.
Corollary 4.7.
Let contact forms $\beta_{1}$ and $\beta_{2}$ on $X$ generate contact structures $\xi_{1}$ and $\xi_{2}$ which belong to the same path-connected component of the space $\mathsf{Cont}(X,Op(\xi^{\partial}))$. Assume that the Reeb vector field $v_{1}$ of the contact form $\beta_{1}$, is traversing.
Then the Reeb vector field $v_{2}$ of a contact form $\beta_{2}$, is traversing as well. As a result, any $v_{2}$-trajectory is homeomorphic to a closed interval or to a singleton (in particular, $v_{2}$ has no closed or trapped trajectories).
Proof.
Consider a smooth path in $\mathsf{Cont}(X,Op(\xi^{\partial}))$ that connects $\xi_{1}$ to $\xi_{2}$. By Lemma 4.7, there exists a smooth orientation-preserving diffeomorphism $\Phi$, identical on $Op(\partial_{1}X)$, such that $\lambda\cdot\beta_{1}=\Phi^{\ast}(\beta_{2})$ for a smooth positive function $\lambda:X\to\mathbb{R}_{+}$. Let us check that $v_{2}=_{\mathsf{def}}\Phi_{\ast}(\lambda^{-1}v_{1})$ is the Reeb vector field for $\beta_{2}$:
$v_{2}\,\rfloor\;d\beta_{2}=\Phi_{\ast}(\lambda^{-1}v_{1})\,\rfloor\;d\beta_{2}=\lambda^{-1}v_{1}\,\rfloor\;\Phi^{\ast}(d\beta_{2})=v_{1}\,\rfloor\;\Big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}+d\beta_{1}\Big{)}=v_{1}\,\rfloor\;\Big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}\Big{)}.$
Since $\big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}\big{)}\big{|}_{\ker(\beta_{1})}=0$, the $1$-form $(v_{2}\,\rfloor\,d\beta_{2})|_{\ker\beta_{1}}=0$. Also, $(v_{2}\,\rfloor\,d\beta_{2})$ vanishes on $v_{1}$:
$v_{1}\,\rfloor(v_{2}\,\rfloor\,d\beta_{2})=v_{1}\,\rfloor\,\big{(}v_{1}\,\rfloor\,\big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}\big{)}\big{)}=0.$
Therefore, $v_{2}\,\rfloor\,d\beta_{2}=0$. Moreover,
$\beta_{2}(v_{2})=\beta_{2}(\Phi_{\ast}(\lambda^{-1}\cdot v_{1}))=\Phi^{\ast}(\beta_{2})(\lambda^{-1}\cdot v_{1})=\lambda\cdot\beta_{1}(\lambda^{-1}\cdot v_{1})=1.$
Hence, $v_{2}=\Phi_{\ast}(\lambda\cdot v_{1})$ is the Reeb vector field for the form $\beta_{2}$. Therefore, the Reeb vector field $v_{2}$ is traversing as well. ∎
We are ready to formulate the main result of this section. It reflects the scheme in (3).
Theorem 4.1.
(Conjugate Holography for Contact Forms and their Reeb Fields)
Let $X_{1},X_{2}$ be compact connected oriented smooth $(2n+1)$-dimensional manifolds with boundaries, equipped with contact forms $\beta_{1},\beta_{2}$. We assume that their Reeb vector fields $v_{1},v_{2}$ are traversing and boundary generic. We denote by $f_{1}^{\bullet},f_{2}^{\bullet}$ the Lyapunov functions such that $df_{1}^{\bullet}(v_{1})=1=df_{2}^{\bullet}(v_{2})$, whose existence is guarantied by Lemma 4.1. In addition, we assume that the $v_{2}$ possesses Property $\mathsf{A}$ from Definition 3.1.222Recall that the property $\partial_{3}X_{2}(v_{2})=\emptyset$ implies Property $\mathsf{A}$.
Let a smooth orientation-preserving diffeomorphism $\Phi^{\partial}:\partial_{1}X_{1}\to\partial_{1}X_{2}$ commute with the two causality maps:
$C_{v_{2}}\circ\Phi^{\partial}=\Phi^{\partial}\circ C_{v_{1}}.$
$\bullet$ Then $\Phi^{\partial}$ extends to a smooth orientation-preserving diffeomorphism $\Phi:X_{1}\to X_{2}$ such that $\Phi$ maps the oriented foliation $\mathcal{F}(v_{1})$ to the oriented foliation $\mathcal{F}(v_{2})$, and the pull-back contact form $\beta^{\dagger}_{1}=\Phi^{\ast}(\beta_{2})$ is $\mathcal{RC}$-equivalent to $\beta_{1}$ (in the sense of Definition 4.5).
$\bullet$ If, in addition, $(\Phi^{\partial})^{\ast}(f_{2}^{\bullet})^{\partial}=(f_{1}^{\bullet})^{\partial}$ and $(\Phi^{\partial})^{\ast}(\beta_{2}|_{\partial_{1}X_{2}})=\beta_{1}|_{\partial_{1}X_{1}},$ then there exists a smooth diffeomorphism $\Phi:X_{1}\to X_{2}$ which extends $\Phi^{\partial}$ and such that $(\Phi)^{\ast}(\beta_{2})=\beta_{1}$.
Proof.
Our arguments rely heavily on the Holography Theorem from [K4]. Assuming Property A, it claims the existence of an orientation-preserving diffeomorphism $\Phi:X_{1}\to X_{2}$ that extends $\Phi^{\partial}$ and maps the $v_{1}$-oriented foliation $\mathcal{F}(v_{1})$ to the $v_{2}$-oriented foliation $\mathcal{F}(v_{2})$.
Let us outline the spirit of the Holography Theorem proof. The reader interested in the technicalities may consult with [K4].
First, using that $v_{2}$ is traversing, we construct a Lyapunov function $f_{2}:X_{2}\to\mathbb{R}$ for it. Then we pull-back, via the diffeomorphism $\Phi^{\partial}$, the restriction $f_{2}^{\partial}=_{\mathsf{def}}f_{2}|_{\partial_{1}X_{2}}$. Since $\Phi^{\partial}$ commutes with the two causality maps, the pull back $f_{1}^{\partial}=_{\mathsf{def}}(\Phi^{\partial})^{\ast}(f^{\partial}_{2})$ has the property $f_{1}^{\partial}(y)>f_{1}^{\partial}(x)$ for any pair $y\succ x$ on the same $v_{1}$-trajectory, the order of points being defined by the $v_{1}$-flow. Equivalently, we get $f_{1}^{\partial}(C_{v_{1}}(x))>f_{1}^{\partial}(x)$ for any $x\in\partial_{1}^{+}X(v_{1})$ such that $C_{v_{1}}(x)\neq x$. As the key step, we prove in [K4] that such $f_{1}^{\partial}$ extends to a smooth function $f_{1}:X_{1}\to\mathbb{R}$ that has the property $df_{1}(v_{1})>0$. Hence, $f_{1}$ is a Lyapunov function for $v_{1}$.
Let $i=1,2$. Recall that each causality map $C_{v_{i}}$ allows to view the $v_{i}$-trajectory space $\mathcal{T}(v_{i})$ as the quotient space $(\partial_{1}X_{i})\big{/}\{C_{v_{i}}(x)\sim x\}$, where $x\in\partial_{1}^{+}X_{i}(v_{i})$; the topology in $\mathcal{T}(v_{i})$ is defined as the quotient topology. Using that $\Phi^{\partial}$ commutes with the causality maps $C_{v_{1}}$ and $C_{v_{2}}$, we conclude that $\Phi^{\partial}$ induces a homeomorphism $\Phi^{\mathcal{T}}:\mathcal{T}(v_{1})\to\mathcal{T}(v_{2})$ of the trajectory spaces, which preserves their natural stratifications.
For a traversing vector field $v_{i}$, the manifold $X_{i}$ carries two mutually transversal foliations: the oriented $1$-dimensional $\mathcal{F}(v_{i})$, generated by the $v_{i}$-flow (or rather by $\hat{v}_{i}$ to insure that the leaves of $\mathcal{F}(\hat{v}_{i})$ are non-singular in the ambient $\hat{X}$), and the foliation $\mathcal{G}(f_{i})$, generated by the constant level hypersurfaces of the Lyapunov function $f_{i}$ (or rather by constant level hypersurfaces of $\hat{f}_{i}:\hat{X}_{i}\to\mathbb{R}$ to insure that the leaves are non-singular in the ambient $\hat{X}$). Note that the leaves of $\mathcal{G}(f_{i})$ may be disconnected, while the leaves of $\mathcal{F}(v_{i})$, the $v_{i}$-trajectories, are connected. The two foliations, $\mathcal{F}(v_{i})$ and $\mathcal{G}(f_{i})$, may serve as a “coordinate grid” on $X_{i}$: every point $x\in X_{i}$ belongs to a unique pair of leaves $\gamma_{x}\subset\mathcal{F}(v_{i})$ and $L_{x}=_{\mathsf{def}}f_{i}^{-1}(f_{i}(x))\subset\mathcal{G}(f_{i})$.
Conversely, using the traversing nature of $v_{i}$, any pair $(y,\,t)$, where $y\in\gamma_{x}\cap\partial_{1}X_{i}$ and $t\in[f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})]\subset\mathbb{R}$ determines a unique point $x\in X_{i}$. Here $[f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})]$ denotes the minimal closed interval that contains the finite set $f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})$. Note that some pairs of leaves $L$ and $\gamma$ may have an empty intersection, and some components of leaves $L$ may have an empty intersection with the boundary $\partial_{1}X_{i}$. However, if $f_{i}(\partial_{1}X_{i})=f_{i}(X_{i})$ as sets, then $f_{i}^{-1}(c)\cap\partial_{1}X_{i}\neq\emptyset$ for any $c\in f_{i}(X_{i})$. In particular, if $\partial_{1}X_{i}$ is connected, then $f_{i}(\partial_{1}X_{i})=f_{i}(X_{i})$ and $f_{i}^{-1}(c)\cap\partial_{1}X_{i}\neq\emptyset$ for any $c\in f_{i}(X_{i})$
In fact, using that $f_{i}$ is a Lyapunov function, the hyprsurface $L=f_{i}^{-1}(c)$ intersects with a trajectory $\gamma$ if and only if $c\in[f_{i}^{\partial}(\gamma\cap\partial_{1}X_{i})]$. Consider a pair of points $y,z\in\partial_{1}X_{i}$. Since the two smooth leaves, $\hat{\gamma}_{y}$ and $\hat{f}_{i}^{-1}(f_{i}(z))$, depend smoothly on the points $y,z$ and are transversal in $\hat{X}_{i}$, their intersection point in $X_{i}$ depends smoothly on $(y,z)\in(\partial_{1}X_{i})\times(\partial_{1}X_{i})$, as long as $f_{i}^{\partial}(z)\in[f_{i}^{\partial}(\gamma_{y}\cap\partial_{1}X_{i})]$. Note that pairs $(y,z)$ with the property $f_{i}^{\partial}(z)\in f_{i}^{\partial}(\gamma_{y}\cap\partial_{1}X_{i})$ give rise to the intersection points that belong to $\partial_{1}X_{i}$.
Now we are ready to extend the diffeomorphism $\Phi^{\partial}$ to a homeomorphism $\Phi:X_{1}\to X_{2}$. In the process, following (3), we assume implicitly the existence of the foliations $\mathcal{F}(v_{i})$ and $\mathcal{G}(f_{i})$ and of the Lyapunov functions $f_{i}$ on $X_{i}$, $i=1,2$.
Take any $x\in X_{1}$. It belongs to a unique pair of leaves $L_{x}\in\mathcal{G}(f_{1})$ and $\gamma_{x}\in\mathcal{F}(v_{1})$. We define $\Phi(x)=x^{\prime}\in X_{2}$, where $x^{\prime}$ is the unique point that belongs to the intersection of $f_{2}^{-1}(f_{1}(x))\in\mathcal{G}(f_{2})$ and the $v_{2}$-trajectory $\gamma^{\prime}=\Gamma_{2}^{-1}(\Phi^{\mathcal{T}}(\gamma_{x}))$. By its construction, $\Phi|_{\partial_{1}X_{1}}=\Phi^{\partial}$. Therefore, $\Phi$ induces the same homeomorphism $\Phi^{\mathcal{T}}:\mathcal{T}(v_{1})\to\mathcal{T}(v_{2})$ as $\Phi^{\partial}$ does.
The leaf-hypersurface $\hat{f}_{2}^{-1}(f_{1}(x))$ depends smoothly on $x$, but the leaf-trajectory $\hat{\gamma}^{\prime}=\Gamma_{2}^{-1}(\Phi^{\mathcal{T}}(\hat{\gamma}_{x}))$ may not! Although the homeomorphism $\Phi$ is a diffeomorphism along the $v_{1}$-trajectories, it is not clear that it is a diffeomorphism on $X_{1}$ (a priori, $\Phi$ is just a Hölder map with a Hölder exponent $\alpha=1/m$, where $m$ is the maximal tangency order of $\gamma$’s to $\partial_{1}X$). Presently, for proving that $\Phi$ is a diffeomorphism, we need Property A [K4]. Assuming its validity, we use the transversality of $\gamma_{x}$ somewhere to $\partial_{1}X$ to claim the smooth dependence of the trajectory $\Gamma_{2}^{-1}(\Phi^{\mathcal{T}}(\hat{\gamma}_{x}))$ on $x$. Since the smooth foliations $\mathcal{F}(\hat{v}_{i})$ and $\mathcal{G}(\hat{f}_{i})$ are transversal, it follows that $x^{\prime}=\Phi(x)$ depends smoothly on $x$. Conjecturally, Property A is unnecessary for establishing that $\Phi$ is a diffeomorphism.
It remains to sort out what happens to the contact forms under this diffeomorphism $\Phi$ or its modifications.
Since $\mathcal{L}_{v_{i}}\beta_{i}=0$, the contact form $\beta_{i}$ is $v_{i}$-invariant, and thus is determined by its boundary data (section) $\beta_{i}^{\partial}:\partial X_{i}\to T^{\ast}(X_{i})|_{\partial X_{i}}$, provided that the $v_{i}$-flow is known (as in the scheme (3)).
We observe that $f_{1}=\Phi^{\ast}(f_{2})$ by the very construction of $\Phi$. Let $v_{1}^{\dagger}=_{\mathsf{def}}\Phi_{\ast}^{-1}(v_{2})$. Since $\Phi$ maps $\mathcal{F}(v_{1})$ to $\mathcal{F}(v_{2})$, we conclude that $v_{1}^{\dagger}$ must be positively-proportional to $v_{1}$. Therefore, for the vector field $v_{1}^{\dagger}$, we get $df_{1}(v_{1}^{\dagger})=\Phi^{\ast}(df_{2})(v_{1}^{\dagger})=df_{2}(v_{2})\;=1$. At the same time, $\beta_{1}(v_{1})=1$.
Consider the $1$-form $\beta_{1}^{\dagger}=_{\mathsf{def}}\Phi^{\ast}(\beta_{2})$ and denote by $\xi_{1}^{\dagger}$ its kernel. Using that $\Phi$ preserves the orientations, we see that $\Phi^{\ast}(\beta_{2})\wedge\Phi^{\ast}(d\beta_{2})^{n}=\Phi^{\ast}(\beta_{2}\wedge(d\beta_{2})^{n})>0$ in $X_{1}$. Therefore, the hyperplane distribution $\xi_{1}^{\dagger}$ is a cooriented contact structure with the contact form $\beta_{1}^{\dagger}$.
By the arguments as in the proof of Corollary 4.7, $v_{1}^{\dagger}=\Phi^{-1}_{\ast}(v_{2})$ is the Reeb vector field for $\beta_{1}^{\dagger}$. Thus, $\beta_{1}^{\dagger}$ is $v_{1}^{\dagger}$-invariant. Since $v_{1}^{\dagger}=\lambda\cdot v_{1}$, where $\lambda:X_{1}\to\mathbb{R}_{+}$ is a positive smooth function, $\beta_{1}$ and $\beta_{1}^{\dagger}$ are $\mathcal{RC}$-equivalent contact forms, and $\xi_{1}$ and $\xi_{1}^{\dagger}$ are $\mathcal{RC}$-equivalent contact structures on $X_{1}$ (see Definition 4.5). Equivalently, $\xi_{2}^{\dagger}=_{\mathsf{def}}\Phi_{\ast}(\xi_{1})$ is $\mathcal{RC}$-equivalent to $\xi_{2}$. This proves the first assertion of the theorem.
Now we turn to the second assertion. The special Lypunov functions $f_{1}^{\bullet},f_{2}^{\bullet}$ are essential for constructing the desired extension $\Phi$ of $\Phi^{\partial}$. We can build the extension $\Phi$ of $\Phi^{\partial}$ as before, using the map $\Phi^{\mathcal{T}}:\mathcal{T}(v_{1})\to\mathcal{T}(v_{2})$ (induced by $\Phi^{\partial}$) and the restrictions $(f_{1}^{\bullet})^{\partial}=(\Phi^{\partial})^{\ast}((f_{2}^{\bullet})^{\partial})$ and $(f_{2}^{\bullet})^{\partial}$ of the Lyapunov functions $f_{1}^{\bullet}$ and $f_{2}^{\bullet}$. By the very construction of $\Phi$, we get $\Phi^{\ast}(f_{2}^{\bullet})=f_{1}^{\bullet}$. Since $df_{i}^{\bullet}(v_{i})=1$ and $\Phi$ maps $v_{1}$-trajectories to $v_{2}$-trajectories, it follows that $\Phi_{\ast}(v_{1})=v_{2}$. By the hypotheses, $(\Phi^{\partial})^{\ast}(\beta_{2}|_{\partial_{1}X_{2}})=\beta_{1}|_{\partial_{1}X_{1}}$ and, since $\beta_{1}(v_{1})=1=\beta_{2}(v_{2})$, we conclude that $(\Phi^{\partial})^{\ast}(\beta_{2}||_{\partial_{1}X_{2}})=\beta_{1}||_{\partial_{1}X_{1}}$, except, perhaps, along the locus $\partial_{2}X_{2}$ where $v_{2}$ is tangent to $\partial_{1}X_{2}$. However, since $\Phi^{\partial}$ is a smooth diffeomorphism, by continuity, we get that $(\Phi^{\partial})^{\ast}(\beta_{2}||_{\partial_{1}X_{2}})=\beta_{1}||_{\partial_{1}X_{1}}$ without exception. Thus, using that $\beta_{i}$ is $v_{i}$-invariant, we get that $(\Phi)^{\ast}(\beta_{2})=\beta_{1}$. ∎
Note that this construction of the extension $\Phi$ from the first assertion of Theorem 4.1 is implicit and far from being canonical. For example, it depends on the choice of extension of $f_{1}^{\partial}=_{\mathsf{def}}(\Phi^{\partial})^{\ast}(f_{2}^{\partial})$ to a smooth function $f_{1}:X_{1}\to\mathbb{R}$, which is strictly monotone along the $v_{1}$-trajectories. The uniqueness (rigidity) of the extension $\Phi$ may be achieved, if we assume the full knowleadge of the manifolds $X_{i}$, equipped with the foliation grids $\mathcal{F}(v_{i}),\mathcal{G}(f_{i})$ and the Lyapunov functions $f_{i}$ or $f_{i}^{\bullet}$ (see Corollary 4.8).
The next theorem, an instant corollary of Theorem 4.1, claims that, assuming Property $\mathsf{A}$, the smooth topological type of the triple $\{X,\,\mathcal{F}(v_{\beta}),\,\mathcal{G}(f^{\bullet})\}$ and of the contact form $\beta$ may be reconstructed from the appropriate boundary-confined data.
Corollary 4.8.
(Reeb Flow Holography)
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold with boundary, equipped with a contact form $\beta$. We assume that the Reeb vector field $v=v_{\beta}$ is traversing and boundary generic and possesses Property $\mathsf{A}$. By Lemma 4.1, using $\beta$, we construct a Lyapunov function $f^{\bullet}:X\to\mathbb{R}$ such that $df^{\bullet}(v)=1$.
Then the following boundary-confined data:
  • the causality map $C_{v}:\partial_{1}^{+}X(v)\to\partial_{1}^{-}X(v)$,
  • the restriction $(f^{\bullet})^{\partial}:\partial_{1}X\to\mathbb{R}$ of the Lyapunov function $f^{\bullet}:X\to\mathbb{R}$,
  • the restriction $\beta^{\partial}:\partial_{1}X\to T^{\ast}X|_{\partial_{1}X}$ of the contact form $\beta$,
are sufficient for a reconstruction of the triple $(X,f^{\bullet},\beta)$ (and thus, of the Reeb vector field $v_{\beta}$ as well), up to diffeomorphisms $\Phi:X\to X$ which are the identity map on $\partial_{1}X$.
Proof.
Assume that we have two sets of structures, $\{X,\,\mathcal{F}(v_{\beta_{1}}),\,\mathcal{G}(f^{\bullet}_{1}),\beta_{1}\}$ and $\{X,\,\mathcal{F}(v_{\beta_{2}}),\,\hfill\break\mathcal{G}(f^{\bullet}_{2}),\beta_{2}\}$, which agree along $\partial_{1}X$. By Theorem 4.1, being applied to $\Phi^{\partial}=\mathsf{id_{\partial_{1}X}}$, the following data
(4.4) $\{\partial_{1}X,\;\mathcal{F}(v_{\beta_{1}})\cap\partial_{1}X,\;\;(f^{\bullet}_{1})^{\partial},\;\beta_{1}^{\partial}\}=\{\partial_{1}X,\;\mathcal{F}(v_{\beta_{2}})\cap\partial_{1}X,\;\;(f^{\bullet}_{2})^{\partial},\;\beta^{\partial}_{2}\}$
along the boundary $\partial_{1}X$ allow an extension of $\Phi^{\partial}$ to a diffeomorphism $\Phi:X\to X$ which maps $\{X,\,\mathcal{F}(v_{\beta_{1}}),\,\mathcal{G}(f^{\bullet}_{1}),\beta_{1}\}$ to $\{X,\,\mathcal{F}(v_{\beta_{2}}),\,\mathcal{G}(f^{\bullet}_{2}),\beta_{2}\}$. Therefore, we have reconstructed the triple $(X,\,f^{\bullet},\,\beta)$, up to a diffeomorphism $\Phi$, which is the identity on $\partial_{1}X$. Note that the reconstruction of the contact form $\beta$ implies the reconstruction of the Reeb field $v_{\beta}$ and of the contact structure $\xi_{\beta}$. ∎
4.1. Boundary data of scattering maps $C_{v_{\beta}}$ and the group-theoretical sections of $\beta$-contactomorphisms
Although the construction of an extension $\Phi:X\to X$ of a diffeomorphism $\Phi^{\partial}:\partial_{1}X\to\partial_{1}X$, subject to the hypotheses of Corollary 4.8, is not canonic, we will argue that by fixing one such extension $\Phi_{\star}$, makes it possible to extend any other diffeomorphism $\tilde{\Phi}^{\partial}:\partial_{1}X\to\partial_{1}X$, subject to the hypotheses of Corollary 4.8, in a rigid way, so that the reconstruction works well for families of such $\tilde{\Phi}^{\partial}$’s.
We assume now that the quadruple $(X,\mathcal{F}(v_{\beta}),f^{\bullet},\beta)$ is known and fixed. We follow the scheme in (3). Then, given any diffeomorphism $\Phi^{\partial}:\partial_{1}X\to\partial_{1}X$ that respects the boundary data (4.4), we can extend it canonically to a diffeomorphism $\Phi:X\to X$ that keeps $\{\mathcal{F}(v),f^{\bullet},\beta\}$ invariant. Indeed, any point $x\in X$ is determined by the unique pair $\big{(}\gamma_{x}\cap\partial_{1}X,\;f^{\bullet}(x)\big{)}$, subject to the requirement $f^{\bullet}(x)\in[f^{\bullet}(\gamma_{x}\cap\partial_{1}X)]$. Now put $\Phi(x)=x^{\prime}$, where the point $x^{\prime}\in X$ is uniquely defined by the pair
$\Big{(}\Phi^{\partial}(\gamma_{x}\cap\partial_{1}X),\;\Phi^{\partial}\big{(}(f^{\bullet})^{-1}(f^{\bullet}(x)\big{)}\;\cap\,\partial_{1}X)\Big{)}.$
By the properties of $\Phi^{\partial}$, the latter pair is of the form
$\big{(}\gamma_{x^{\prime}}\cap\partial_{1}X,\;\big{(}(f^{\bullet})^{-1}(f^{\bullet}(x^{\prime})\big{)}\cap\partial_{1}X)\big{)},$
where $f^{\bullet}(x^{\prime})\in f^{\bullet}([\gamma_{x^{\prime}}\cap\partial_{1}X])$.
Assuming the validity of Property $\mathsf{A}$, as in Theorem 4.1, we get that $\Phi$ is a smooth diffeomorphism.
Since $\Phi^{\partial}$ preserves the boundary data $\{(f^{\bullet})^{\partial},\beta^{\partial}\}$ and the contact form $\beta$ is $v_{\beta}$-invariant, we get that $\Phi$ must preserve both $\beta$ and its Reeb vector field $v_{\beta}$. Thus, the extension $\Phi$ is a contactomorphism.
Consider the topological groups $\mathsf{Diff}^{\infty}_{+}(X)$ and $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ of smooth orientation-preserving diffeomorphisms and the natural restriction homomorphism $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(X)\to\mathsf{Diff}^{\infty}_{+}(\partial_{1}X).$ Similarly, we may consider the groups ${}^{\circ}\mathsf{Diff}^{\infty}_{+}(X)$ and ${}^{\circ}\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ that are the connected components of $\mathsf{Diff}^{\infty}_{+}(X)$ and $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ that contain the identity.
Recall that the famous Cerf Theorem [Cerf] claims that any orientation-preserving diffeomorphism of $S^{3}$ is isotopic to the identity, and thus, extends to a diffeomorphism of the $4$-ball $D^{4}$. Therefore, the homomorphism $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(D^{4})\to\mathsf{Diff}^{\infty}_{+}(S^{3})$ is onto.
However, remarkably, for any smooth compact connected $X$ with boundary, the restriction map $\mathcal{R}:\;^{\circ}\mathsf{Diff}^{\infty}_{+}(X)\to\;^{\circ}\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ does not admit a group-theoretical section ([ChMa], Corollary 1.8)!
Let $\mathsf{Diff}^{\infty}_{+}(X;\beta,f^{\bullet})$ denote the subgroup of the group $\mathsf{Diff}^{\infty}_{+}(X)$ of smooth orientation-preserving diffeomorphisms, whose elements preserve the contact form $\beta$ and the Lyapunov function $f^{\bullet}:X\to\mathbb{R}$ such that $df^{\bullet}(v_{\beta})=1$.
Let $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X;\,C_{v_{\beta}},\beta^{\partial},(f^{\bullet})^{\partial})$ denote the subgroup of $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ whose elements commute with the causality map $C_{v_{\beta}}:\partial_{1}^{+}X(v_{\beta})\to\partial_{1}^{-}X(v_{\beta})$, preserve the restriction to the boundary $(f^{\bullet})^{\partial}:\partial_{1}X\to\mathbb{R}$ of the Lyapunov function $f^{\bullet}$ and the section $\beta^{\partial}:\partial_{1}X\to T^{\ast}(X)|_{\partial_{1}X}$.
In contrast with the fact that $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(X)\to\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ does not admit a group-theoretical section [ChMa], we get the following claim.
Theorem 4.2.
Under the hypotheses of Corollary 4.8, the restriction homomorphism $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(X)\to\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ admits a group-theoretical continuous section $\sigma$ over the subgroup $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X;\,C_{v_{\beta}},\beta^{\partial},(f^{\bullet})^{\partial})\subset\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$.
Moreover, the image $\sigma(\mathsf{Diff}^{\infty}_{+}(\partial_{1}X;\,C_{v_{\beta}},\beta^{\partial},(f^{\bullet})^{\partial}))$ belongs to the subgroup $\mathsf{Diff}^{\infty}_{+}(X;\beta,f^{\bullet})$ of $\mathsf{Diff}^{\infty}_{+}(X)$.
Proof.
Follows instantly from the previous arguments and the proof of Theorem 4.1. ∎
5. On contact forms and contact vector fields on manifolds with boundary
For the reader convenience, we present a few basic properties of contact forms and contact vector fields we will rely upon.
Let $\alpha$ be a differential $k$-form on a smooth manifold $X$ and let $Z\subset X$ be a smooth submanifold. We use somewhat unconventional notations: we write “$\alpha|_{Z}$” for the restriction of the section $\alpha:X\to\bigwedge^{k}T^{\ast}X$ to the section $\alpha|:Z\to\bigwedge^{k}T^{\ast}Z$, and we write “$\alpha||_{Z}$” for the restriction of $\alpha$ to the section of the bundle $\bigwedge^{k}T^{\ast}X|_{Z}\to Z$.
Let $\alpha^{\partial}=_{\mathsf{def}}\;\alpha||_{\partial_{1}X}$. In particular, $\beta^{\partial}=_{\mathsf{def}}\beta||_{\partial_{1}X}$ is the restriction of the section-form $\beta$ to the section of the bundle $T^{\ast}X\to\partial_{1}X$ (not to be confused with the restriction of $\beta$ to the submanifold $\partial_{1}X\subset X$).
Definition 5.1.
A smooth vector field $u$ on $X$ is called contact field for a given contact $1$-form $\beta$ if $\mathcal{L}_{u}\beta=\lambda\cdot\beta$ for some smooth function $\lambda:X\to\mathbb{R}$. $\diamondsuit$
Since $\mathcal{L}_{u}(\mu\beta)=(d\mu(u)+\mu\lambda)\beta$, where $\lambda,\mu$ are smooth functions and $\mathcal{L}_{u}(\beta)=\lambda\beta$, the notion of contact vector field depends only on the contact structure $\xi_{\beta}$, generated by a contact form $\beta$.
To derive Corollary 5.2 of Theorem 4.1 we will rely on a couple of lemmas.
Lemma 5.1.
Let $v$ be a traversing vector field on a compact connected oriented smooth $(2n+1)$-dimensional manifold $X$ with boundary. Assume that two cooriented contact structures $\xi_{1}$ and $\xi_{2}$ on $X$ share the same Reeb vector field $v$ of their contact forms $\beta_{1}$ and $\beta_{2}$. [^3] Hence, $\xi_{1}$ and $\xi_{2}$ are $\mathcal{R}$-equivalent.
If the distributions-subbundles $\xi_{1}^{\partial}\subset T_{\ast}(X)\,|_{\partial_{1}X}\text{ and \,}\xi_{2}^{\partial}\subset T_{\ast}(X)\,|_{\partial_{1}X}$ coincide, then $\xi_{1}=\xi_{2}$.
Proof.
Let $\beta_{i}$ ($i=1,2$) be a $1$-form whose kernel is the hyperplane distribution $\xi_{i}$. By the definition of Reeb vector field $v$, we have $v\rfloor d\beta_{i}=0$ and $v\rfloor\beta_{i}=1$. The Cartan’s identity $\mathcal{L}_{v}\beta_{i}=v\rfloor d\beta_{i}+d(v\rfloor\beta_{i})$ implies that the directional derivative $\mathcal{L}_{v}\beta_{i}=0$, i.e., $\beta_{i}$ is $v$-invariant. Thus, the $1$-form $\beta_{i}^{\partial}$, viewed as a section of the bundle $T^{\ast}X|_{\partial_{1}^{+}X(v)}$, spreads by the $v$-flow uniquely along each $v$-trajectory to produce the invariant $1$-form $\beta_{i}$. Using that $v$ is traversing, we conclude that $\beta_{i}^{\partial}$ determines $\beta_{i}$ everywhere in $X$. As a result, the distributions $\xi_{1},\xi_{2}$ are $v$-invariant. Thus, $\xi_{1}$ is determined by $\xi_{1}^{\partial}$ and $\xi_{2}$ is determined by $\xi_{2}^{\partial}$. Therefore, if $\xi_{1}^{\partial}=\xi_{2}^{\partial}$, then $\xi_{1}=\xi_{2}$. ∎
Given a contact form $\beta$, any vector field on $X$ can be written as $u=h\cdot v_{\beta}+w$, where $v_{\beta}$ is the Reeb vector field, $h:X\to\mathbb{R}$ is a smooth function, and the vector field $w\in\xi_{\beta}$. Then $\mathcal{L}_{u}\beta=\lambda\cdot\beta$ for some smooth function $\lambda:X\to\mathbb{R}$ if and only if
$\mathcal{L}_{u}\beta\equiv\mathcal{L}_{h\cdot v_{\beta}+w}\beta=h\cdot\mathcal{L}_{v_{\beta}}\beta+(v_{\beta}\,\rfloor\,\beta)\,dh+\mathcal{L}_{w}\beta=\lambda\cdot\beta.$
Using that $\mathcal{L}_{v_{\beta}}\beta\equiv 0$ and $v_{\beta}\,\rfloor\,\beta=1$, this is equivalent to the condition $1\cdot dh+\mathcal{L}_{w}\beta=\lambda\cdot\beta,$ or, by Cartan’s formula, to the condition
(5.1) $w\,\rfloor\,d\beta=-dh+\lambda\cdot\beta.$
Since $v_{\beta}\,\rfloor\,d\beta=\mathbf{0}$, (5.1) implies that
$0=-w\,\rfloor\,(v_{\beta}\,\rfloor\,d\beta)=v_{\beta}\,\rfloor\,(w\,\rfloor\,d\beta)=v_{\beta}\,\rfloor\,(-dh+\lambda\cdot\beta)=-dh(v_{\beta})+\lambda.$
Thus, $\lambda=dh(v_{\beta})$, and the equation (5.1) transforms into the equation
(5.2) $w\,\rfloor\,d\beta=-dh+dh(v_{\beta})\cdot\beta$
with respect to $w\in\xi_{\beta}$. Restricting (5.2) to the hypersurface $\{h=c\}$ gives the equations
(5.3) $(w\,\rfloor\,d\beta)\,|_{h^{-1}(c)}$ $=$ $-dh(v_{\beta})\cdot\beta\,|_{h^{-1}(c)},$
$(w\,\rfloor\,\beta)\,|_{h^{-1}(c)}$ $=$ $0$
Lemma 5.2.
Let $\beta$ be a contact form. For any choice of the smooth function $h$, there exists a unique vector field $w\in\ker\beta$ that satisfies the equation (5.2).
Proof.
Since $d\beta|_{\ker(\beta)}$ is a non-degenerated skew-symmetric bilinear form, for any $1$-form $\alpha$, viewed as a section of the dual bundle $(\ker(\beta))^{\ast}\to X$, there exists a section $w(\alpha)$ of the bundle $\ker(\beta)\to X$ such that $w(\alpha)\,\rfloor\,d\beta=\alpha$. Therefore, the restriction of equation (5.2) to $\ker(\beta)$ gives the solvable equation
(5.4) $(w\,\rfloor\,d\beta)|_{\ker(\beta)}=-(dh)|_{\ker(\beta)}$
with respect to the vector field $w\in\ker\beta$. By the same token, the solution $w$ of (5.4) is unique.
If $w\in\ker\beta$ is the solution of (5.4), then it is automatically a solution of (5.2), since $v_{\beta}\,\rfloor\,(w\,\rfloor\,d\beta)=v_{\beta}\,\rfloor\,(-dh+dh(v_{\beta}))\cdot\beta$ produces the identity $0=0$. ∎
Lemma 5.3.
Let $\beta$ be a contact form. For any choice of smooth function $h$, the vector field $w\in\ker\beta$ that satisfies (5.2) is tangent to any hypersurface $\{h=c\}_{c\in\mathbb{R}}$.
Proof.
It follows from (5.2) that $0=w\,\rfloor\,(w\,\rfloor\,d\beta)=w\,\rfloor\,(-dh+dh(v_{\beta})\cdot\beta)$. Since $w\,\rfloor\,\beta=0$, we get $w\,\rfloor\,dh=0$. Thus, $w$ is tangent to any hypersurface $h^{-1}(c)\subset X$, where $c\in\mathbb{R}$. ∎
By Lemma 5.3, $u=h\cdot v_{\beta}+w$ is also tangent to $\partial_{1}X$, since $h|_{\partial_{1}X}=0$. Thus, we getting the following claim.
Corollary 5.1.
Let $\beta$ be a contact form on $X$. For any smooth function $h:\hat{X}\to\mathbb{R}$ which has the properties listed in (4), and for any vector field $w\in\ker(\beta)$, which satisfies the equation
$w\,\rfloor\,d\beta=-dh+dh(v_{\beta})\cdot\beta,$
the one-parameter family diffeomorphisms $\{\phi_{t}(w):X\to X\}_{t}$ that integrates the field $w$ is well-defined for all $t\in\mathbb{R}$.
Also, the one-parameter family diffeomorphisms $\{\psi_{t}(u):X\to X\}_{t}$ that integrates the contact vector field $u=h\cdot v_{\beta}+w$, which satisfies the equation $\mathcal{L}_{u}\beta=dh(v_{\beta})\cdot\beta,$ is well-defined for all $t\in\mathbb{R}$. $\diamondsuit$
Lemma 5.4.
For a vector field $w\in\ker(\beta)$, $\mathcal{L}_{w}\beta=0$ if and only if $w=0$.
Proof.
By the formula $\mathcal{L}_{w}\beta=w\rfloor d\beta+d(w\rfloor\beta)=w\rfloor d\beta$, we get $\mathcal{L}_{w}\beta=0$ if and only if $w\rfloor d\beta=0$. Since the symplectic form $d\beta|_{\ker\beta}$ is non-degenerated, it follows that $w=0$. ∎
The next example exhibits the simplest cases of these contact forms and fields.
Example 5.1.
$\bullet$ Let $X=\{x^{2}+y^{2}+z^{2}\leq 1\}\subset\mathbb{R}^{3}$ and $\beta=dz+x\,dy$. The Reeb vector field $v_{\beta}$ is $\partial_{z}$. It is tangent to the $2$-sphere $\partial_{1}X$ along its equator $E=\{z=0\}\cap X$. Thus, the restriction of $\beta$ to $E$ is equal to the restriction of the form $x\,dy$ to $E$. If $\theta$ is the natural parameter on the circle $E$, then $\beta|_{E}=cos^{2}(\theta)\,d\theta$. It is non-negative everywhere on $E$ and vanishes at $\theta=\pm\pi/2$, or at $(x,y,z)=(0,\pm 1,0)$.
If in (5.2), we pick $h=(x^{2}+y^{2}+z^{2}-1)/2$ then the equation (5.2) reduces to
$w\,\rfloor\,(dx\wedge dy)=-(x\,dx+y\,dy+z\,dz)+z(dz+x\,dy).$
Coupling this equation with the equation $w\,\rfloor\,(dz+x\,dy)=0$, we get a formula
$w=(xz-y)\,\partial_{x}+x\,\partial_{y}-x^{2}\,\partial_{z}\;$
for the vector field $w\in\xi_{\beta}$ which is tangent to the concentric spheres. Note that the fixed point set of the $w$-flow is the $z$-axis. Finally, put
$u=h\cdot v_{\beta}+w=(xz-y)\partial_{x}+x\,\partial_{y}+\frac{1}{2}\big{(}-x^{2}+y^{2}+z^{2}-1\big{)}\partial_{z}.$
The $u$-flow preserves the distribution $\xi_{\beta}$ and the sphere $\partial_{1}X$. It has exactly at two fixed points $(0,0,\pm 1)$.
$\bullet$ We pick the shell $\{1\leq x^{2}+y^{2}+z^{2}\leq 4\}$ for the role of $X$ and the same contact form $\beta=dz+x\,dy$. Then $\partial_{2}^{+}X(v_{\beta})$ is the equator $E_{1}=\{x^{2}+y^{2}+z^{2}=1\}\cap\{z=0\}$ and $\partial_{2}^{-}X(v_{\beta})$ is the equator $E_{2}=\{x^{2}+y^{2}+z^{2}=4\}\cap\{z=0\}$. The same computation leads to $\beta|_{E_{1}}\geq 0$ and $\beta|_{E_{2}}\leq 0$. In other words, ±$\pm\beta|_{\partial_{2}^{\pm}X(v_{\beta})}\geq 0$.
Now the function $h=(x^{2}+y^{2}+z^{2}-4)(x^{2}+y^{2}+z^{2}-1)$. We leave to the reader to solve the equation $w\,\rfloor\,d\beta=-dh+dh(v_{\beta})\cdot\beta$ for $w\in\xi_{\beta}$.
$\bullet$ The next example is also given by $\beta=dz+x\,dy$, but $h=z$. The choice of $X\subset\mathbb{R}^{3}$ is unimportant. In this case, $v_{\beta}=\partial_{z}$ and (5.2) reduces to
$w\,\rfloor\,(dx\wedge dy)=-dz+(dz+x\,dy)=x\,dy.$
Thus, $w=x\,\partial_{x}\;\in\xi_{\beta}$ and $u=x\,\partial_{x}+z\partial_{z}$. The vector field $w$ is tangent to the planes $\{z=c\}_{c}$ and the vector field $u$ is tangent to the plane $\{z=0\}$. The $u$-flow keeps the form $\beta$ invariant. The plane $\{x=0\}$ is the fixed point set of the $w$-flow and the line $\{x=0,z=0\}$ is the fixed point set of the $u$-flow. $\diamondsuit$
Example 5.2.
$\bullet$ Let $X=\{x_{1}^{2}+y_{1}^{2}+x_{2}^{2}+y_{2}^{2}+z^{2}\leq 1\}\subset\mathbb{R}^{5}$ and $\beta=dz+x_{1}dy_{1}+x_{2}dy_{2}$. We pick $h=(x_{1}^{2}+y_{1}^{2}+x_{2}^{2}+y_{2}^{2}+z^{2}-1)/2$. The Reeb vector field $v_{\beta}=\partial_{z}$ and the $3$-form
$\beta\wedge d\beta=dz\wedge dx_{1}\wedge dy_{1}+dz\wedge dx_{2}\wedge dy_{2}+x_{1}\,dy_{1}\wedge dx_{2}\wedge dy_{2}+x_{2}\,dy_{2}\wedge dx_{1}\wedge dy_{1}.$
Let $E$ be the equator $\{z=0\}\cap\{x_{1}^{2}+y_{1}^{2}+x_{2}^{2}+y_{2}^{2}=1\}$, a $3$-sphere $\partial_{2}^{-}X(v_{\beta})\subset\partial_{1}X=S^{4}$. Thus $\beta\wedge d\beta|_{E}=(x_{1}\,dy_{1}\wedge dx_{2}\wedge dy_{2}+x_{2}\,dy_{2}\wedge dx_{1}\wedge dy_{1})|_{E}$.
A somewhat lengthy computation in the spherical coordinates $\psi,\theta\in[0,\pi],\phi\in[0,2\pi]$ on the $3$-sphere $E$ leads to the formula
$\beta\wedge d\beta|_{E}=\big{[}(sin^{2}\psi\,cos^{2}\,\psi+sin^{4}\psi\,sin^{2}\theta\,cos^{2}\phi)\,sin\theta\big{]}\;d\psi\wedge d\theta\wedge d\phi\;\geq 0.$
[…]
These examples motivate the following conjecture.
Conjecture 5.1.
Let $\beta$ be a smooth contact form on a compact connected manifold $X$ and $v_{\beta}$ its boundary generic Reeb vector field. Then:
  • For an even $j$,  $\beta|_{\mathsf{int}(\partial_{j}^{+}X(v_{\beta}))}$ is a contact form away from a zero measure set. Moreover,
    $\mp\big{[}\beta\wedge(d\beta)^{\frac{2n-j}{2}}\big{]}\big{|}_{\partial_{j}^{\pm}X(v_{\beta})}\geq 0,\text{ and }\beta\wedge(d\beta)^{\frac{2n-j}{2}}\big{|}\big{|}_{\partial_{j+1}X(v_{\beta})}\equiv 0\;.$
  • For an odd $j$, $d\beta|_{\mathsf{int}(\partial_{j}^{+}X(v_{\beta}))}$ is a symplectic form away from a zero measure set. Moreover,
    ±$\pm(d\beta)^{\frac{2n+1-j}{2}}\big{|}_{\partial_{j}^{\pm}X(v_{\beta})}\geq 0,\text{ and }(d\beta)^{\frac{2n+1-j}{2}}\big{|}\big{|}_{\partial_{j+1}X(v_{\beta})}\equiv 0\;.$
    $\diamondsuit$
Recall that the Gray Stability Theorem [Gray] claims that, for any smooth family $\{\xi_{t}\}_{t\in[0,1]}$ of contact structures on a closed smooth manifold $X$, there exists a smooth isotopy $\{\phi_{t}\}_{t\in[0,1]}$ such that $(\phi_{t})^{\ast}(\xi_{t})=\xi_{0}$ for all $t\in[0,1]$. In fact, we may assume that $X$ is compact in this claim, provided $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}$, where $\dot{\beta}_{t}=_{\mathsf{def}}\frac{d}{dt}\beta_{t}$.
Lemma 5.5.
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold $X$ with boundary. Let $\{h_{t}:\hat{X}\to\mathbb{R}\}_{t\in[0,1]}$ be a family of smooth functions, each of which has 00 as its regular value, $h_{t}^{-1}(0)=\partial_{1}X$, and $h_{t}^{-1}((-\infty,0])=X$.
Consider a family $\{\xi_{t}\}_{t\in[0,1]}$ of cooriented contact structures on $X$, generated by a smooth family of contact forms $\{\beta_{t}\}_{t\in[0,1]}$, such that the $t$-derivative $\dot{\beta}_{t}=dh_{t}$ along $\partial_{1}X$ [^4] The property $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}$ implies that $\beta_{t}|_{\partial_{1}X}$ is constant.
Then there is a smooth $t$-family of diffeomorphism $\{\phi_{t}:X\to X\}_{t\in[0,1]}$, such that $(\phi_{t})^{\ast}(\xi_{t})=\xi_{0}$; in particular, $\phi_{1}$ is a contactomorphism for the pair $\xi_{0},\;\xi_{1}$.
Proof.
We apply the Moser trick [Mo], slightly adjusted to the category of manifolds with boundary. We are looking for a $t$-family $\phi_{t}:X\to X$ of diffeomorphisms such that $(\phi_{t})^{\ast}(\beta_{t})=\lambda_{t}\cdot\beta_{0}$, where $\lambda_{t}:X\to\mathbb{R}_{+}$ is a smooth $t$-family of positive functions. Let $\{w_{t}\}$ be a $t$-family of vector fields on $X$ to be constructed, such that the $w_{t}$-flow gives rise to the desired family of diffeomorphisms $\phi_{t}:X\to X$.
We use [Ge], Lemma 2.19, which claims that
$\frac{d}{dt}(\phi_{t}^{\ast}(\beta_{t}))=\phi_{t}^{\ast}(\dot{\beta}_{t}+\mathcal{L}_{w_{t}}\beta_{t}).$
By differentiating the identity $\phi_{t}^{\ast}(\beta_{t})=\lambda_{t}\cdot\beta_{0}$ with respect to $t$, the desired property of the vector fields $w_{t}$ can be stated as
$\phi_{t}^{\ast}(\dot{\beta}_{t}+\mathcal{L}_{w_{t}}\beta_{t})\,=\,\dot{\lambda}_{t}\cdot\beta_{0}\,=\,\frac{\dot{\lambda}_{t}}{\lambda_{t}}\cdot\phi_{t}^{\ast}(\beta_{t}).$
By introducing the auxiliary function
$\mu_{t}=_{\mathsf{def}}\,\frac{d}{dt}\big{(}log(\lambda_{t})\big{)}\circ(\phi_{t})^{-1},$
with the help of Cartan’s formula, we get
(5.5) $\phi_{t}^{\ast}\big{(}\dot{\beta}_{t}+d(\beta_{t}(w_{t}))+w_{t}\,\rfloor\,d\beta_{t}\big{)}=\phi_{t}^{\ast}(\mu_{t}\cdot\beta_{t}).$
If we choose $w_{t}\in\xi_{t}=\ker(\beta_{t})$, this equation will be satisfied if $\dot{\beta}_{t}+w_{t}\rfloor d\beta_{t}=\mu_{t}\cdot\beta_{t}$, or equivalently, if
(5.6) $w_{t}\,\rfloor\,d\beta_{t}=\mu_{t}\cdot\beta_{t}-\dot{\beta}_{t}.$
Plugging the Reeb vector field $v_{\beta_{t}}$ of $\beta_{t}$ in (5.6), we get
(5.7) $\dot{\beta}_{t}(v_{\beta_{t}})=\mu_{t}.$
Hence, we can use (5.7) to determine the function $\mu_{t}$.
With $\mu_{t}$ in place, the non-degeneracy of the $2$-form $d\beta_{t}|_{\xi_{t}}$ and the fact that (thanks to (5.7)) $v_{\beta_{t}}\in\ker(\mu_{t}\cdot\beta_{t}-\dot{\beta}_{t})$ allows us to find a unique solution $w_{t}\in\xi_{t}$ of (5.6), written as
(5.8) $w_{t}\,\rfloor\,d\beta_{t}=\dot{\beta}_{t}(v_{\beta_{t}})\cdot\beta_{t}-\dot{\beta}_{t}.$
It remains to verify that this $w_{t}$ is tangent to the boundary $\partial_{1}X$, provided that $\dot{\beta}_{t}=dh_{t}$ along $\partial_{1}X$. Applying the operation $w_{t}\,\rfloor\,$ to both sides of (5.8), we get $0=\dot{\beta}_{t}(w_{t})$. By the hypothesis, $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}$. Therefore, $0=\dot{\beta}_{t}(w_{t})||_{\partial_{1}X}=dh_{t}(w_{t})||_{\partial_{1}X}$. It follows that $dh_{t}(w_{t})|_{\partial_{1}X}=dh_{t}(w_{t})|_{h_{t}^{-1}(0)}=0$, which implies that $w_{t}$ is tangent to $h_{t}^{-1}(0)=\partial_{1}X$. Therefore, the $w_{t}$-flow preserves the manifold $X$ and the diffeomorphisms $\phi_{t}:X\to X$ are well-defined for all $t\in[0,1]$. Thus, $\phi_{1}^{\ast}(\beta_{1})=\lambda_{1}\cdot\beta_{0}$ and $\phi_{1}^{\ast}(\xi_{1})=\xi_{0}$. ∎
Corollary 5.2.
We adopt the notations and hypotheses of Lemma 5.5: let a $t$-family of contact forms $\{\beta_{t}\}_{t\in[0,1]}$ on $X$ be such that the $t$-derivative $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}.$ Let $\{\xi_{t}\}_{t\in[0,1]}$ be the corresponding $t$-family of contact structures.
Assume that the Reeb vector field $v_{0}$ of the contact form $\beta_{0}$, is traversing.
Then the Reeb vector field $v_{1}$ of a contact form $\beta_{1}$, is traversing as well. As a result, any $v_{1}$-trajectory is homeomorphic to a closed interval or to a singleton (in particular, $v_{1}$ has no closed or trapped trajectories).
If, in addition, $v_{0}$ is boundary generic, then the causality maps $C_{v_{1}}$ and $C_{v_{0}}$ are conjugate with the help of the diffeomorphism $\phi_{1}^{\partial}=\phi_{1}|:\partial_{1}X\to\partial_{1}X$ as in Lemma 5.5.
Moreover, the knowledge of the boundary data $\{C_{v_{\beta_{0}}},(f^{\bullet}_{0})^{\partial},\beta_{0}^{\partial}\}$ is sufficient for reconstructing the smooth topological type of $\{X,\xi_{1}\}$, provided that $v_{\beta_{0}}$ satisfies Property $\mathsf{A}$.
Proof.
By Lemma 5.5, for an appropriate diffeomorphism $\phi_{1}:X\to X$, we get $\phi_{1}^{\ast}(\beta_{1})=\lambda_{1}\cdot\beta_{0}$. By arguments as in the proof of Corollary 4.7, we get that $v_{1}=(\phi_{1})_{\ast}(\lambda_{1}^{-1}v_{0})$ is the Reeb vector field for $\beta_{1}$. This implies that, with the help of $\phi_{1}:X\to X$, the smooth dynamics of the $v_{0}$-flow and the $v_{1}$-flow are similar. In particular, if $v_{0}$ is traversing, then so is $v_{1}$. If $v_{0}$ is boundary generic, then so is $v_{1}$. For the same reason, the causality maps $C_{v_{1}}$ and $C_{v_{0}}$, which depend only on the conformal classes of vector fields $v_{0}$ and $v_{1}$, are conjugate with the help of $\phi_{1}^{\partial}$.
The last assertion follows by Combining Theorem 4.1 (or Corollary 4.8) with Lemma 5.5, provided that $v_{\beta_{0}}$ satisfies Property $\mathsf{A}$. ∎
6. Non-squeezing of $2n$-symplectic volumes, Morse wrinkles of Reeb fields, and holography
Let us introduce now some new numerical invariants of contact forms on manifolds with boundary.
Definition 6.1.
Let $v_{\beta}$ be a traversing Reeb vector field for a contact $1$-form $\beta$ on $X$. The Reeb diameter of $(X,\beta)$ is defined by
$\mathsf{diam}_{\mathcal{R}}(\beta)\;=_{\mathsf{def}}\;\sup_{\gamma\subset X}\Big{\{}\int_{\gamma}\beta\Big{\}},$
where $\gamma$ runs over the $v_{\beta}$-trajectories. $\diamondsuit$
Lemma 6.1.
Consider an orientation-preserving submersion $\Psi:X\to Y$ of two compact equidimentional manifolds. Let $\beta_{Y}$ be a contact form on $Y$ whose Reeb vector field $v_{Y}$ is traversing. We assume that $\Psi(\partial_{1}X)$ is boundary generic relative to $v_{Y}$.
Then the form $\beta_{X}=_{\mathsf{def}}\Psi^{\ast}(\beta_{Y})$ is a contact form on $X$. Its Reeb vector field $v_{X}$ is traversing with the help of the Lyapunov function $f_{X}=\Psi^{\ast}(f_{Y})$, where $df_{Y}(v_{Y})>0$. Moreover, the Reeb diameters satisfy the inequality
$\mathsf{diam}_{\mathcal{R}}(\beta_{X})\leq\mathsf{diam}_{\mathcal{R}}(\beta_{Y}).$
If $\Psi$ is a covering map (i.e., an immersion that is also surjective), then
$\mathsf{diam}_{\mathcal{R}}(\beta_{X})=\mathsf{diam}_{\mathcal{R}}(\beta_{Y}).$
Proof.
Since $\Psi$ is submersion and $\Psi(\partial_{1}X)$ is boundary generic relative to $v_{Y}$, for any $v_{Y}$-trajectory $\gamma$ (which is a closed interval or a singleton) the set $\Psi^{-1}(\gamma)$ is either a disjoint union of finitely many closed intervals $\tilde{\gamma}$, or a finite set. Moreover, $\Psi:\tilde{\gamma}\hookrightarrow\gamma$ is an immersion. Thus, $\int_{\tilde{\gamma}}\beta_{X}=\int_{\tilde{\gamma}}\Psi^{\ast}(\beta_{Y})=\int_{\Psi(\tilde{\gamma})}\beta_{Y}\leq\int_{\gamma}\beta_{Y}$. In the case of a covering $\Psi$, we get $\Psi:\tilde{\gamma}\to\gamma$ is a diffeomorphism. In such a case, $\int_{\tilde{\gamma}}\beta_{X}=\int_{\gamma}\beta_{Y}$. Since any $v_{X}$-trajectory $\tilde{\gamma}$ is obtained from some $v_{Y}$-trajectory $\gamma$ as described above, the claims about the Reeb diameters follow. ∎
Definition 6.2.
For a traversing Reeb field, we introduce the volume of the space of Reeb trajectories $\mathcal{T}(v_{\beta})$ by the formula
(6.1) $vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta}))=_{\mathsf{def}}\;\int_{\partial_{1}^{+}X(v)}(d\beta)^{n}.$
$\diamondsuit$
Among the contact forms $\beta$ that represent a given contact structure $\xi$, we may consider only $\beta$’s such that their Reeb vector fields $v_{\beta}$ are traversing and $\mathsf{diam}_{\mathcal{R}}(\beta)=1$. Of course, if $\mathsf{diam}_{\mathcal{R}}(\beta)=\lambda>0$ for some contact form $\beta$ and its traversing Reeb vector field $v_{\beta}$, then $\mathsf{diam}_{\mathcal{R}}(\lambda^{-1}\beta)=1$ for the traversing Reeb vector field $\lambda v$ of the contact form $\lambda^{-1}\beta$. Thus, without lost of generality, we may assume that $\mathsf{diam}_{\mathcal{R}}(\beta)=1$ in Theorem 6.1.
Theorem 6.1.
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold with boundary, equipped with a contact $1$-form $\beta$. Let $v_{\beta}$ be a traversing boundary generic Reeb vector field for $\beta$.
Then we get an “isoperimetric” inequality of volumes
(6.2) $\int_{X}\beta\wedge(d\beta)^{n}\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\int_{\partial_{1}^{+}X(v_{\beta})}(d\beta)^{n}=_{\mathsf{def}}\;\mathsf{diam}_{\mathcal{R}}(\beta)\cdot vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta})).$
This enequality can be also written in the “equatorial” form:
(6.3) $\int_{X}\beta\wedge(d\beta)^{n}\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\int_{\partial_{2}X(v_{\beta})}\beta\wedge(d\beta)^{n-1}.$
Proof.
The argument is a variation on the theme of Santalo formula [S]. For any $y\in\partial_{1}^{+}X(v_{\beta})$ we consider the $v$-trajectory $\gamma_{y}$ through $y$. The $(-v)$-directed projection $p:X\to\partial_{1}^{+}X(v_{\beta})$ is a submersion, away from the set $p^{-1}(\partial_{2}X(v_{\beta}))$ of zero measure in $\partial_{1}^{+}X(v_{\beta})$. Thus, we can use the Dieudonne generalization of the Fubini formula [D]:
$\int_{X}\beta\wedge(d\beta)^{n}=\int_{y\in\partial_{1}^{+}X(v_{\beta})}\Big{(}\int_{\gamma_{y}}\beta\Big{)}(d\beta)^{n}\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\int_{\partial_{1}^{+}X(v_{\beta})}(d\beta)^{n},$
together with Definition 6.1, to validate (6.2). By Stokes’ theorem, applied to $\partial_{1}^{+}X(v_{\beta})$, (6.3) follows.
Although this argument is simple, it is somewhat surprising to realize the $(d\beta)^{n}$-volume of the trajectory space $\mathcal{T}(v_{\beta})$ depends only on the form $\beta\wedge(\partial\beta)^{n-1}|_{\partial_{2}X(v_{\beta})}$ on the “equator” $\partial_{2}X(v_{\beta})$. ∎
Definition 6.3.
Assume that the Reeb vector field $v_{\beta}$ of a contact form $\beta$ is traversing. We introduce the average length of the Reeb trajectories by
(6.4) $\mathsf{av}_{\mathcal{R}}(\beta)\;=_{\mathsf{def}}\;\frac{\int_{X}\beta\wedge(d\beta)^{n}}{\int_{\partial_{1}^{+}X(v_{\beta})}(d\beta)^{n}}.$
By (6.3), $\mathsf{av}_{\mathcal{R}}(\beta)\;\leq\;\mathsf{diam}_{\mathcal{R}}(\beta).$diamondsuit
Example 7.1. Let $\beta=dz+x\,dy$ and $X=\big{\{}\frac{x^{2}}{(A/2)^{2}}+\frac{y^{2}}{(B/2)^{2}}+\frac{z^{2}}{(C/2)^{2}}\leq 1\big{\}}$. Then the inequality (6.2) can be written as
$\frac{\pi}{6}ABC\leq C\cdot\big{(}\frac{\pi}{4}AB\big{)},$
or as $\frac{\pi}{6}\leq\frac{\pi}{4}$. $\diamondsuit$
Corollary 6.1.
For any $\eta\in C^{\infty}(\mathcal{T}(v_{\beta}))$, the contact form $\beta_{1}=\beta+d\eta$ shares with $\beta$ the Reeb diameter $\mathsf{diam}_{\mathcal{R}}(\beta)$, the volume $\int_{X}\beta\wedge(d\beta)^{n}$, and the volume $vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta}))$ of the trajectory space. $\diamondsuit$
Proof.
The “exact slantings” $\beta_{1}=\beta+d\eta$ of a contact form $\beta$ do not change the $\beta$-lengths of Reeb trajectories, provided the function $\eta\in C^{\infty}(\mathcal{T}(v_{\beta}))$.
By Lemma 4.5, $\beta$ and $\beta+d\eta$ share the same Reeb vector field: $v_{\beta}=v_{\beta+d\eta}$. Thus, for a trajectory $\gamma$ bounded by points $a$ and $b$, we get
$\int_{\gamma}(\beta+d\eta)=\int_{\gamma}\beta+(\eta(b)-\eta(a))=\int_{\gamma}\beta$
since $\eta(b)=\eta(a)$. As a result, $\mathsf{diam}_{\mathcal{R}}(\beta+d\eta)=\mathsf{diam}_{\mathcal{R}}(\beta)$.
Using the Fubini formula, we verify that the $(\beta+d\eta)\wedge(d(\beta+d\eta))^{n}$-induced and the $(\beta)\wedge(d\beta)^{n}$-induced volumes of $X$ are equal:
$\int_{X}(\beta+d\eta)\wedge(d(\beta+d\eta))^{n}=\int_{y\in\partial_{1}^{+}X(v_{\beta})}\Big{(}\int_{\gamma_{y}}(\beta+d\eta)\Big{)}(d\beta)^{n}$
[…]
$\int_{y\in\partial_{1}^{+}X(v_{\beta})}\Big{(}\int_{\gamma_{y}}\beta\Big{)}(d\beta)^{n}=\int_{X}(\beta)\wedge(d\beta)^{n}.$
Also,
$vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta}))=\int_{\partial_{1}^{+}X(v_{\beta})}(d\beta)^{n}=\int_{\partial_{1}^{+}X(v_{\beta+d\eta})}(d\beta)^{n}$
$=\int_{\partial_{1}^{+}X(v_{\beta+d\eta})}(d(\beta+d\eta))^{n}=vol_{(d(\beta+d\eta))^{n}}(\mathcal{T}(v_{\beta+d\eta}))$
Definition 6.4.
Let $Y$ be a smooth manifold and $v$ a non-vanishing vector field on it that admits a Lyapunov function. Consider a compact smooth connected manifold $X$ with boundary, $\dim X=\dim Y$, and a regular imbedding $\Psi:X\to\mathsf{int}(Y)$ such that $v$ is boundary generic relative the hypersurface $\Psi(\partial_{1}X)$. We introduce a natural number $c^{\bullet}(\Psi,v)$ as half of the maximum number of times a $v$-trajectory intersects $\Psi(\partial_{1}X)$. $\diamondsuit$
In general, $c^{\bullet}(\Psi,v)$ resists an intrinsic characterization in terms of the pull-back $\Psi^{\dagger}(v)$ of the vector field $v$ to $X$. We know only one lower bound of $c^{\bullet}(\Psi,v)$ in terms of $\Psi^{\dagger}(v)$:
$c^{\bullet}(\Psi^{\dagger}(v))=_{\mathsf{def}}\max_{\gamma^{\dagger}\subset X}\Big{\{}\frac{1}{2}|\omega(\gamma^{\dagger})|\Big{\}}\leq c^{\bullet}(\Psi,v),$
where $\gamma^{\dagger}$ is a $\Psi^{\dagger}(v)$-trajectory (see (2.3)). This estimate follows from the local models of the traversing boundary generic flows in the vicinity of their trajectories [K1], [K3].
Remark 6.1.
Even if the connected source and the target of $\Psi$ both carry convex vector fields, the number $c^{\bullet}(\Psi,v)$ can be arbitrary big. Indeed, just rap a very narrow ($\epsilon\ll 1$) band $Z=I\times[-\epsilon,\epsilon]$ with rounded corners $k$ times around the cylinder $C=S^{1}\times[-1,1]$, where $I$ denotes the interval $[0,1]$ and the vector field $v$ is aligned with the generators of the cylinder. Next, consider the embedding
$\Psi:X=S^{1}\times I\times[-\epsilon,\epsilon]\;\hookrightarrow\;S^{1}\times S^{1}\times[-1,1]=Y,$
obtained by multiplying the wrapping map $\psi:Z\subset C$ with the identity map on the first multiplayer $S^{1}$. Then $c^{\bullet}(\Psi,v_{Y})=k$. $\diamondsuit$
Consider a traversing $v$-flow on a compact manifold $Y$ and a regular embedding $\Psi:X\to\mathsf{int}(Y)$, where $X$ is a compact manifold of the same dimension as $Y$. Assuming that $v$ is boundary generic with respect to the hypersurface $\Psi(\partial_{1}X)$, with the help of $(-v)$-flow, we can generate a shadow $\mathcal{X}^{\partial}(\Psi)$ of $\Psi(X)$ on the boundary of $Y$. It is produced by forming the space $\mathcal{X}(\Psi)$ of $v$-trajectories through the points of $\Psi(X)$ (equivalently, through the locus $\partial_{1}^{+}\Psi(X)(v)$) and then taking the intersection $\mathcal{X}(\Psi)\cap\partial_{1}^{+}Y(v)$. This constructions presumes that $\Psi(X)$ is “semi-transparent” for the flow (see Fig. 1, the left diagram).
The next theorem resembles Gromov’s non-squeezing theorem [Gr2] in symplectic geometry, being transferred to the contact geometry environment (see Fig. 1). It deals with the embeddings $\Psi:X_{1}\to X_{2}$ ($\dim X_{1}=2n+1=\dim X_{2}$) that respect the contact forms: i.e., $\beta_{1}=\Psi^{\ast}(\beta_{2})$. It employs the number $c^{\bullet}(\Psi,v_{\beta_{2}})$ that measures how folded is the image $\Psi(\partial_{1}X_{1})$ with respest to the $v_{\beta_{2}}$-flow. The “equatorial” inequality of Theorem 6.2 may be also viewed as delivering a lower bound of the number $c^{\bullet}(\Psi,v_{\beta_{2}})$ in terms of the appropriate $\beta_{1}$- and $\beta_{2}$-generated symplectic $2n$-volumes.
Theorem 6.2.
(Non-squeezing equatorial property of the Reeb fields)
Consider two compact smooth manifolds $X_{1},X_{2}$, equipped with contact forms $\beta_{1},\beta_{2}$, respectively. We assume that the Reeb vector fields $v_{\beta_{1}},v_{\beta_{2}}$ are boundary generic and $v_{\beta_{2}}$ is traversing. Let $\Psi:X_{1}\hookrightarrow X_{2}$ be a smooth embedding such that $\beta_{1}=\Psi^{\ast}(\beta_{2})$. Then
$\int_{X_{1}}\beta_{1}\wedge(d\beta_{1})^{n}\;\leq\;\int_{X_{2}}\beta_{2}\wedge(d\beta_{2})^{n},$
$\mathsf{diam}_{\mathcal{R}}(\beta_{1})\;\leq\;\mathsf{diam}_{\mathcal{R}}(\beta_{2}).$
We also get the inequality between the symplectic $2n$-volumes of the two trajectory spaces (or, equivalently, between the contact-like volumes of the tangency locii, the “equators”):
$0\leq\int_{\partial_{2}X_{1}(v_{\beta_{1}})}\beta_{1}\wedge(d\beta_{1})^{n-1}=vol_{(d\beta_{1})^{n}}(\mathcal{T}(v_{\beta_{1}}))\;\leq\;$
$\;\leq\;c^{\bullet}(\Psi,v_{\beta_{2}})\cdot vol_{(d\beta_{2})^{n}}(\mathcal{T}(v_{\beta_{2}}))=c^{\bullet}(\Psi,v_{\beta_{2}})\cdot\int_{\partial_{2}X_{2}(v_{\beta_{2}})}\beta_{2}\wedge(d\beta_{2})^{n-1}.$
Figure 1. The figure helps with the notations in the proof of Theorem 6.2. In the figure, $v_{\beta_{1}}$ and $v_{\beta_{2}}$ are the constant vertical vector fields and $c^{\bullet}(\Psi,v_{\beta_{2}})=2$.
Proof.
To simplify the notations, put $v_{1}=v_{\beta_{1}}$, $v_{2}=v_{\beta_{2}}$.
Using that $\beta_{1}=\Psi^{\ast}(\beta_{2})$, we see that $\Psi_{\ast}(v_{1})=v_{2}|_{\Psi(X_{1})}$. Moreover, the combinatorial tangency pattern of every $v_{1}$-trajectory $\gamma_{1}$ to $\partial_{1}X_{1}$ coincides with the tangency pattern of the portion $\Psi(\gamma_{1})$ of the unique $v_{2}$-trajectory $\gamma_{2}\supset\Psi(\gamma_{1})$ to $\Psi(\partial_{1}X)$. Note that the tangency pattern of $\gamma_{2}$ to $\Psi(\partial_{1}X)$ could be different.
The validation of the first two inequalities is straightforward. Since $\Psi^{\ast}(\beta_{2})=\beta_{1}$, by naturality, we get $\int_{X_{1}}\beta_{1}\wedge(d\beta_{1})^{n}=\int_{\Psi(X_{1})}\beta_{2}\wedge(d\beta_{2})^{n}$. Evidently, $\int_{\Psi(X_{1})}\beta_{2}\wedge(d\beta_{2})^{n}\leq\int_{X_{2}}\beta_{2}\wedge(d\beta_{2})^{n}$. Thus, $\int_{X_{1}}\beta_{1}\wedge(d\beta_{1})^{n}\leq\int_{X_{2}}\beta_{2}\wedge(d\beta_{2})^{n}$.
It follows instantly that $\mathsf{diam}_{\mathcal{R}}(\beta_{1})\leq\mathsf{diam}_{\mathcal{R}}(\beta_{2})$, since the $\Psi$-image of any $v_{\beta_{1}}$-trajectory $\gamma$ is a segment of a $v_{\beta_{2}}$ trajectory $\gamma^{\prime}$ that contains $\Psi(\gamma)$.
Now we are going to validate the equatorial inequality (see Fig. 1). Let $\mathcal{W}_{2}(\Psi)$ denote the $2n$-dimensional set comprised of $v_{2}$-trajectories through the locus $\Psi(\partial_{2}X_{1}(v_{1}))\coprod\partial_{2}X_{2}(v_{2})$. Consider the set (see Fig.  1)
$\mathcal{V}_{1}^{\circ}(\Psi)=_{\mathsf{def}}\Psi\big{(}\partial_{1}^{+}X_{1}(v_{1})\big{)}\setminus\big{(}\Psi(\partial_{1}^{+}X_{1}(v_{1}))\cap\mathcal{W}_{2}(\Psi)\big{)}$
along which the continuous dependence of the $v_{2}$-flow on the initial conditions in $\Psi(\partial_{1}^{+}X_{1}(v_{1}))$ holds. Let $\mathcal{V}^{\circ}_{1,\kappa}(\Psi)$ be a connected component of the set $\mathcal{V}_{1}^{\circ}(\Psi)$ and let $\mathcal{V}_{1,\kappa}(\Psi)$ be the closure of $\mathcal{V}_{1,\kappa}^{\circ}(\Psi)$. We denote by $\partial^{+}_{1}X^{\circ}_{1,\kappa}(v_{1})\subset\partial_{1}^{+}X(v_{1})$ the $\Psi$-preimage of $\mathcal{V}^{\circ}_{1,\kappa}(\Psi)$.
Keeping in mind that $\Psi_{\ast}(v_{1})=v_{2}|_{\Psi(X_{1})}$, consider the set $\mathcal{X}_{\kappa}^{\circ}(\Psi)$, formed by the union of portions of the $(-v_{2})$-trajectories that originate at the set $\mathcal{V}_{1,\kappa}^{\circ}(\Psi)$ and terminate at the points of the set $\partial_{1}^{+}X_{2}(v_{2})$. Let $\mathcal{X}_{\kappa}(\Psi)$ be the closure of $\mathcal{X}_{\kappa}^{\circ}(\Psi)$.
Let $\mathcal{Y}_{\kappa}(\Psi)=_{\mathsf{def}}\mathcal{X}_{\kappa}(\Psi)\cap\partial_{1}^{+}X_{2}(v_{2})$ be the set where the $v_{2}$-trajectories through $\mathcal{V}_{1,\kappa}(\Psi)$ terminate. Then the boundary of the set $\mathcal{X}_{\kappa}(\Psi)$ consists of three parts: $\mathcal{V}_{1,\kappa}(\Psi)),\,\mathcal{Y}_{\kappa}(\Psi)$, and the rest, denoted by $\mathcal{Z}_{\kappa}(\Psi)$. The $2n$-dimensional Lebesgue measures of mutual intersections of these three parts is zero.
We notice that $\mathcal{Z}_{\kappa}(\Psi)$ consists of portions of $(-v_{2})$-trajectories. Since $v_{2}\,\rfloor\,d\beta_{2}=0$, we conclude that $\int_{\mathcal{Z}_{\kappa}(\Psi)}(d\beta_{2})^{n}=0$.
Now we apply Stokes’ theorem to the zero form $d\big{(}(d\beta_{2})^{n}\big{)}$ on $\mathcal{X}_{\kappa}(\Psi)$ to get
$0=\int_{\mathcal{X}_{\kappa}(\Psi)}d((d\beta_{2})^{n})=\int_{\mathcal{Y}_{\kappa}(\Psi)}(d\beta_{2})^{n}-\int_{\mathcal{V}_{1,\kappa}(\Psi)}(d\beta_{2})^{n}.$
Therefore, using that $\Psi(\partial^{+}_{1}X^{\circ}_{1,\kappa}(v_{1}))=\mathcal{V}_{1,\kappa}(\Psi)$ by the definition of the relevant sets, we get
$\int_{\mathcal{Y}_{\kappa}(\Psi)}(d\beta_{2})^{n}=\int_{\mathcal{V}_{1,\kappa}(\Psi)}(d\beta_{2})^{n}=\int_{\Psi(\partial^{+}_{1}X^{\circ}_{1,\kappa}(v_{1}))}(d\beta_{2})^{n}.$
Thus, by the latter equality,
$\int_{\partial^{+}_{1}X^{\circ}_{1,\kappa}(v_{1})}(d\beta_{1})^{n}=\int_{\Psi(\partial^{+}_{1}X^{\circ}_{1,\kappa}(v_{1}))}(d\beta_{2})^{n}=\int_{\mathcal{Y}_{\kappa}(\Psi)}(d\beta_{2})^{n}\leq\int_{\partial_{1}^{+}X_{2}(v_{2}))}(d\beta_{2})^{n},$
since $\mathcal{Y}_{\kappa}(\Psi)$ is a part of the locus $\partial_{1}^{+}X_{2}(v_{2}))$ on which, by Lemma 4.3, $(d\beta_{2})^{n}\geq 0$.
Next, we notice that the domains $\{\mathcal{Y}_{\kappa}(\Psi)\}_{\kappa}$ may overlap massively in $\partial_{1}^{+}X_{2}(v_{2}))$ and the multiplicity of the overlap is bounded from above by the maximal number of times a $v_{2}$-trajectory, passing through the set $\Psi(X_{1})$, hits the set $\Psi(\partial_{1}^{+}X_{1}(v_{1}))$. By Definition  6.4, this natural number is $c^{\bullet}(\Psi,v_{\beta_{2}})$.
Adding the inequalities for all $\kappa$ and using that $\{\partial_{1}^{+}X^{\circ}_{1,\kappa}(v_{1})\}_{\kappa}$ form a partition of $\partial_{1}^{+}X_{1}(v_{1})$, modulo the set of $2n$-dimensional measure zero, we get
$\int_{\partial_{1}^{+}X_{1}(v_{1})}(d\beta_{1})^{n}=\sum_{\kappa}\int_{\partial_{1}^{+}X_{1,\kappa}(v_{1})}(d\beta_{1})^{n}\,\leq\,c^{\bullet}(\Psi,v_{\beta_{2}})\cdot\int_{\partial_{1}^{+}X_{2}(v_{2}))}(d\beta_{2})^{n}.$
Finally, in view of Definition 6.2 and by Stokes’ Theorem, we get
$\int_{\partial_{2}X_{1}(v_{1})}\beta_{1}\wedge(d\beta_{1})^{n-1}=vol_{(d\beta_{1})^{n}}(\mathcal{T}(v_{1})),\;\;\int_{\partial_{2}X_{1}(v_{2})}\beta_{2}\wedge(d\beta_{2})^{n-1}=vol_{(d\beta_{2})^{n}}(\mathcal{T}(v_{2})).$
Example 6.1.
Let $Y$ be a compact connected smooth $(2n+1)$-manifold with boundary, equipped with a contact form $\beta_{Y}$ such that its Reeb vector field $v_{Y}$ is traversing and boundary generic. Let $X$ be the ellipsoid $\big{\{}\sum_{i=1}^{2n+1}\frac{x_{i}^{2}}{a_{i}^{2}}\leq 1\big{\}}$ in the Euclidean space $\mathsf{E}^{2n+1}$, equipped with the standard contact form
$\beta=dx_{2n+1}+\frac{1}{2}\sum_{i=1}^{n}(x_{i+1}\,dx_{i}-x_{i}\,dx_{i+1}).$
Then Theorem 6.2 claims that, for any embedding $\Psi:X\to Y$ such that $\beta=\Psi^{\ast}(\beta_{Y})$, the following inequalities are valid:
  • $\frac{\pi^{n+\frac{1}{2}}}{\Gamma(n+\frac{3}{2})}\;a_{1}a_{2}\dots a_{2n+1}\leq\int_{Y}\beta_{Y}\wedge(d\beta_{Y})^{n},$
  • $a_{2n+1}\leq\mathsf{diam}_{\mathcal{R}}(\beta_{Y}),$
  • $\frac{\pi^{n}}{\Gamma(n+1)}\;a_{1}a_{2}\dots a_{2n}\leq c^{\bullet}(\Psi,v_{\beta_{2}})\cdot\int_{\partial_{1}^{+}Y(v_{Y})}(d\beta_{Y})^{n}$
[…]
where $\Gamma(\sim)$ is the Gamma function. In particular, for $X$ being a ball of radius $r$, we get the following constraints on the injectivity radius $r$ of a contact embedding:
$\frac{\pi^{n+\frac{1}{2}}}{\Gamma(n+\frac{3}{2})}\;r^{2n+1}\leq\int_{Y}\beta_{Y}\wedge(d\beta_{Y})^{n},$
$r\leq\mathsf{diam}_{\mathcal{R}}(\beta_{Y}),$
[…]
$\diamondsuit$
Example 6.2.
Consider a solid $X$ shaped as a sand clock (see Fig. 2) which is vertically alined with the $z$-axis in the $xyz$-space $\mathbb{R}^{3}$. Let $\beta=dz+x\,dy$. We consider a $2$-parameter family of transformations $\{A_{\lambda\mu}:x\to\lambda x,y\to\lambda y,z\to\mu z\}_{\lambda,\mu\in\mathbb{R}_{+}}$. We compare two contact $3$-folds $(X,\beta)$ and $(A_{\lambda\mu}(X),\beta)$. Then $c^{\bullet}(\Psi,v_{2})\geq 2$.
For $c^{\bullet}(\Psi,v_{2})=2$, unless $\lambda<2,\,\mu<1,\,\lambda^{2}\mu<1$, no contact embedding $\Psi$ of $X$ into $A_{\lambda\mu}(X)$ exists. $\diamondsuit$
Let $v_{\beta}$ be a boundary generic Reeb vector field of a contact form $\beta$ on a compact $(2n+1)$-manifold $X$. For any boundary generic $v_{\beta}$, $\partial_{j}X(v_{\beta})=\partial(\partial_{j-1}^{+}X(v_{\beta}))$ for all $j\geq 1$. Thus, for an odd $j$, by Stokes’ theorem,
[…]
In contrast, for an even $j$, we get
$\int_{\partial_{j}X(v_{\beta})}\beta\wedge(d\beta)^{\frac{2n-j}{2}}=\int_{\partial_{j-1}^{+}X(v_{\beta})}d\big{(}\beta\wedge(d\beta)^{\frac{2n-j}{2}}\big{)}=\int_{\partial_{j-1}^{+}X(v_{\beta})}(d\beta)^{\frac{2n+2-j}{2}}.$
Conjecturally, the last integral is positive (see Conjecture 5.1 and Examples 5.1-5.2).
Motivated by these observations, we introduce some numerical measures of the stratified concavity/convexity of the Reeb flows $v_{\beta}$ relative to the boundary $\partial_{1}X$:
$\kappa_{j}^{+}(\beta)$ $=_{\mathsf{def}}\;$ $\int_{\partial_{j}^{+}X(v_{\beta})}(d\beta)^{\frac{2n+1-j}{2}}\;\;\;(=\kappa_{j+1}(\beta)),$
[…]
$\kappa_{j}^{+}(\beta)$ $=_{\mathsf{def}}\;$ $\int_{\partial_{j}^{+}X(v_{\beta})}\beta\wedge(d\beta)^{\frac{2n-j}{2}},$
(6.6) $\kappa_{j}(\beta)$ $=_{\mathsf{def}}\;$ $\int_{\partial_{j}X(v)}\beta\wedge(d\beta)^{\frac{2n-j}{2}}\;\;\;(=\kappa^{+}_{j-1}(\beta))\text{\; for an even }j\in[2,2n],$
Of course, these quantities depend on the choice of the contact form $\beta$ that represents a given contact structure $\xi$ and whose Reeb vector field $v_{\beta}$ is boundary generic. They measure how “wrinkled” $\partial_{1}X$ is in relation to the Reeb flow $v_{\beta}$. By the very definition of $\kappa_{j}^{+}(\beta),\kappa_{j}(\beta)$, they belong to the category of “boundary data”.
Note that the inequality (6.2) can be written as $\kappa^{+}_{0}(\beta)\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\kappa^{+}_{1}(\beta)$, and the inequality (6.3) as $\kappa^{+}_{0}(\beta)\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\kappa_{2}(\beta)$, where $\kappa^{+}_{0}(\beta)=_{\mathsf{def}}\int_{X}\beta\wedge(d\beta)^{n}$.
6.1. Invariance of the volumes $\{\kappa_{j}(\beta)\}$ under special deformations of $\beta$
Definition 6.5.
Let =$\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ be a family of contact forms on a compact connected smooth $(2n+1)$-manifold $X$ with boundary.
$\bullet$ We say the family $\mathcal{B}$ is variationally boundary exact if
$\frac{\partial}{\partial t}\beta_{t}\Big{|}_{\partial_{1}X}=d\eta_{t}$
for a family of smooth functions $\eta_{t}:\partial_{1}X\to\mathbb{R}$.
$\bullet$ We say the family $\mathcal{B}$ is variationally boundary $(2n+1-j)$-exact if, for some smooth family of $(2n-j)$-forms $\{\zeta_{t}\}_{t}$ on $\partial_{1}X$,
(6.7) $\Big{\{}\frac{\partial}{\partial t}\Big{[}\beta_{t}\wedge(d\beta_{t})^{\frac{2n-j}{2}}\Big{]}\Big{\}}\Big{|}_{\partial_{1}X}=d\zeta_{t},\text{ where }j\in[2,2n]\text{ is even, }$,
(6.8) $\Big{\{}\frac{\partial}{\partial t}\Big{[}(d\beta_{t})^{\frac{2n+1-j}{2}}\Big{]}\Big{\}}\Big{|}_{\partial_{1}X}=d\zeta_{t},\text{ where }j\in[1,2n-1]\text{ is odd }.$
Occasionally, we compliment these requirements with the additional one:
(6.9) $\zeta_{t}\big{|}_{\partial_{j+1}X(v_{\beta_{t}})}=0.\quad\quad\quad\quad\quad\quad\quad\quad\hfill\diamondsuit$
Lemma 6.2.
Any variationally boundary exact family $\mathcal{B}$ of contact forms is variationally boundary $(2n+1-j)$-exact.
If the functions ${\eta_{t}}$ are such that $\eta_{t}|_{\partial_{j+1}X(v_{t})}=0$, then $\zeta_{t}|_{\partial_{j+1}X(v_{t})}=0$.
Proof.
Note that, for any $t$-family differential forms, their exterior differential $d=_{\mathsf{def}}d_{t}$ commutes with the partial derivative operator $\frac{\partial}{\partial t}$ and with the restriction of forms to $\partial_{1}X$. Consider the derivative
$\frac{\partial}{\partial t}\Big{[}(d\beta_{t})^{k}\Big{]}=k\;\frac{\partial}{\partial t}\Big{(}d\beta_{t}\Big{)}\wedge\big{[}(d\beta_{t})^{k-1}\big{]}=k\;d\Big{(}\frac{\partial}{\partial t}(\beta_{t})\Big{)}\wedge\big{[}(d\beta_{t})^{k-1}\big{]}.$
Let $\eta_{t}:\partial_{1}X\to\mathbb{R}$ be a smooth family of functions from Definition 6.5. Replacing $\frac{\partial}{\partial t}(\beta_{t})\big{|}_{\partial_{1}X}$ with $d\eta_{t}$ in these formulas and using that $d(d\eta_{t})=0$, we see that $\frac{\partial}{\partial t}\Big{[}(d\beta_{t})^{k}\Big{]}\Big{|}_{\partial_{1}X}=0$. Similarly,
$\frac{\partial}{\partial t}\Big{[}\beta_{t}\wedge(d\beta_{t})^{k}\Big{]}=\frac{\partial}{\partial t}\Big{[}\beta_{t}\Big{]}\wedge(d\beta_{t})^{k}+\beta_{t}\wedge\frac{\partial}{\partial t}\Big{[}(d\beta_{t})^{k}\Big{]}$
Since $\frac{\partial}{\partial t}\Big{[}(d\beta_{t})^{k}\Big{]}\Big{|}_{\partial_{1}X}=0$, this formula collapses to $\frac{\partial}{\partial t}\Big{[}\beta_{t}\Big{]}\wedge(d\beta_{t})^{k}\Big{|}_{\partial_{1}X}$.
Replacing $\frac{\partial}{\partial t}\Big{[}\beta_{t}\Big{]}\Big{|}_{\partial_{1}X}$ with $d\eta_{t}$ once more, we get
$\frac{\partial}{\partial t}\Big{[}\beta_{t}\Big{]}\wedge(d\beta_{t})^{k}\Big{|}_{\partial_{1}X}=d\eta_{t}\wedge(d\beta_{t})^{k}\Big{|}_{\partial_{1}X}=d\Big{[}\eta_{t}\cdot(d\beta_{t})^{k}\Big{]}\Big{|}_{\partial_{1}X}=d\zeta_{t}.$
Thus, (6.7) is valid. If the functions ${\eta_{t}}$ are such that $\eta_{t}|_{\partial_{j+1}X(v_{t})}=0$, then $\zeta_{t}|_{\partial_{j+1}X(v_{t})}=0$; so the property (6.9) holds as well. ∎
Proposition 6.1.
Let =$\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ be a variationally boundary $(2n+1-j)$-exact family of contact forms on a compact connected smooth $(2n+1)$-manifold $X$. Assume that their Reeb vector fields $v_{\beta_{t}}$ are boundary generic for $t=0,1$, and for almost all $t\in(0,1)$.
$\bullet$ Then the integral $\kappa_{j}(\beta_{t})$ ($j\equiv 0\mod 2$) from (6.6) and $\kappa_{j-1}^{+}(\beta_{t})$ ($j\equiv 0\mod 2$) from (6) are $t$-independent. If, in addition, the property (6.9) holds, then $\kappa^{+}_{j}(\beta_{t})$ ($j\equiv 0\mod 2$) is also invariant within $\mathcal{B}$.
$\bullet$ If =$\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ is variationally boundary exact family, then $\{\kappa_{j}(\beta_{t})\}$ are $\mathcal{B}$-invariant for all even $j\in[2,2n]$. If in addition, (6.9) holds for the function $\eta_{t}$, then $\kappa^{+}_{j}(\beta_{t})$ ($j\equiv 0\mod 2$) is $\mathcal{B}$-invariant.
Proof.
Let $\mathsf{d}$ be the exterior derivative operator on $X\times[0,1]$ and $d_{t}$ the exterior derivative operator on a typical slice $X\times\{t\}$. Let $v_{t}=_{\mathsf{def}}v_{\beta_{t}}$. Consider the following sets:
$\mathcal{Y}_{j}=_{\mathsf{def}}\bigcup_{t\in[0,1]}\partial_{j}X(v_{t})\times\{t\}\subset X\times[0,1],$
$\mathcal{Y}_{j}^{+}=_{\mathsf{def}}\bigcup_{t\in[0,1]}\partial_{j}^{+}X(v_{t})\times\{t\}\subset X\times[0,1],$
Although $\mathcal{Y}_{j}$ may have singularities, located in some $t$-slices, by the hypothesis, such special $t$’s have measure zero in $[0,1]$; so one can integrate a smooth differential form against the rectifiable chain $\mathcal{Y}_{j}$.
We start with an even $j$. Put $s=\frac{2n-j}{2}$. Then
$\kappa_{j}(\beta_{1})-\kappa_{j}(\beta_{0})=_{\mathsf{def}}\int_{\partial_{j}X(v_{1})}\beta_{1}\wedge(d_{1}\beta_{1})^{s}-\int_{\partial_{j}X(v_{0})}\beta_{0}\wedge(d_{0}\beta_{0})^{s}\stackrel{{\scriptstyle Stokes}}{{=}}\quad\int_{\mathcal{Y}_{j}}\mathsf{d}\big{(}\beta_{t}\wedge(d_{t}\beta_{t})^{s}\big{)}$
$=\int_{\mathcal{Y}_{j}}(d_{t}\beta_{t})^{s+1}\;+\int_{\mathcal{Y}_{j}}\frac{\partial}{\partial t}\Big{[}\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big{]}\wedge dt$
Each of these two integrals vanishes. Indeed, $dt$ is not present in the $(2s+2)$-form $(d_{t}\beta_{t})^{s+1}$. Thus its restriction to the $(2s+2)$-chain $\mathcal{Y}_{j}\subset\partial_{1}X\times[0,1]$ vanishes. The second integral is also zero by the variationally boundary $(2n+1-j)$-exact hypothesis:
[…]
since integrating the exact form $d_{t}\zeta_{t}$ over the cycle $\partial_{j}X(v_{\beta_{t}})\subset\partial_{1}X$ results in zero. Therefore, $\kappa_{j}(\beta_{1})=\kappa_{j}(\beta_{0})$ for this particular even $j\in[2,2n]$.
By Lemma 6.2, the second bullet’s claim follows for all even $j\in[2,2n]$.
Using Stokes’ theorem identity $\kappa_{j}(\beta)=\kappa_{j-1}^{+}(\beta)$ ($j\equiv 0\mod 2$), we conclude that, for odd $j\in[1,2n-1]$, the quantities $\{\kappa_{j}^{+}(\beta)\}_{j}$ are $\mathcal{B}$-deformation invariants as well.
If we add property (6.9) to property (6.7), then a similar computation shows that $\kappa^{+}_{j}(\beta_{1})=\kappa^{+}_{j}(\beta_{0})$ for an even $j$. This time, Stokes’ theorem is applied to $\mathcal{Y}_{j}^{+}$. Its boundary consists of tree parts $\partial_{j}^{+}X(v_{0}),\partial_{j}^{+}X(v_{1})$, and their complement $\delta\mathcal{Y}_{j}^{+}=\mathcal{Y}_{j+1}$. Therefore,
$\kappa^{+}_{j}(\beta_{1})-\kappa^{+}_{j}(\beta_{0})=\int_{\mathcal{Y}^{+}_{j}}\mathsf{d}\big{(}\beta_{t}\wedge(d_{t}\beta_{t})^{s}\big{)}\pm\int_{\delta\mathcal{Y}^{+}_{j}}\beta_{t}\wedge(d_{t}\beta_{t})^{s}$
$=\int_{\mathcal{Y}^{+}_{j}}(d_{t}\beta_{t})^{s+1}\;+\int_{\mathcal{Y}^{+}_{j}}\frac{\partial}{\partial t}\Big{[}\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big{]}\wedge dt\;\pm\;\int_{\delta\mathcal{Y}^{+}_{j}}\beta_{t}\wedge(d_{t}\beta_{t})^{s}.$
Since $dt$ is not present in the first and third integrals, they vanish. As a result,
[…]
by (6.9). ∎
Consider now the following quantities (see Definition 6.3):
[…]
They measure how “wrinkled” the Morse stratification $\{\partial_{j}^{\pm}X(v_{\beta})\}_{j}$ could be with respect to a given contact structure $\xi=\ker\beta$ on $X$. As a practical matter, the computation of $\{\mathcal{K}_{j}^{+}(\xi)\}_{j},\{\mathcal{K}_{j}(\xi)\}_{j}$ could be quite challenging…
[…]
Theorem 6.3.
(Holography for the $\kappa_{j}^{+}$-volumes under the contact embeddings)
Let $X$ and $Y$ be compact smooth $(2n+1)$-manifolds, equipped with contact forms $\beta_{X},\beta_{Y}$ and their Reeb vector fields $v_{X}$, $v_{Y}$, respectively. We assume that $v_{X}$ and $v_{Y}$ are boundary generic and $v_{Y}$ is traversing. Let $\Psi:X\hookrightarrow\mathsf{int}(Y)$ be a smooth embedding such that $\beta_{X}=\Psi^{\ast}(\beta_{Y})$. For $j\in[1,2n+1]$, consider the $(-v_{Y})$-guided projections $\Pi_{Y}:\Psi\big{(}\partial_{j}^{+}X(v_{X})\big{)}\to\partial_{1}^{+}Y(v_{Y})$ and denote their images by $\mathcal{Y}^{+}_{j}$. [^5] Here, as before, we assume that $\Psi(X)$ is “semi-transparent” for the $(-v_{Y})$-flow.
Then the integrals $\{\kappa_{j}^{+}(\beta_{X})\}_{j}$ from (6) and $\{\kappa_{j}(\beta_{X})\}_{j}$ from (6.6) can be recovered from the $\Pi_{Y}$-projections $\{\mathcal{Y}^{+}_{j}\}_{j}$ and from the knowledge of the contact form $\beta_{Y}^{\partial_{+}}=_{\mathsf{def}}\beta_{Y}||_{\partial_{1}^{+}Y(v_{Y})}$. Specifically, for an odd $j\in[1,2n-1]$,
$\kappa_{j}^{+}(\beta_{X})=-\int_{\mathcal{Y}_{j}^{+}}(d\beta_{Y})^{\frac{2n+1-j}{2}}$
and, for an even $j\in[2,2n]$,
$\kappa_{j}(\beta_{X})=-\int_{\mathcal{Y}_{j-1}^{+}}(d\beta_{Y})^{\frac{2n+2-j}{2}}.$
Proof.
We consider the union $\mathcal{X}_{j}^{+}$ of downward $(-v_{Y})$-directed trajectories through the points of $\Psi(\partial^{+}_{j}(X)(v_{X}))$. Note that $\dim(\mathcal{X}_{j}^{+})=2n+2-j$. Let
$\mathcal{Y}_{j}^{+}=_{\mathsf{def}}\mathcal{X}_{j}^{+}\cap\partial_{1}^{+}Y(v_{Y})$
—the $\Pi_{Y}$-image of $\partial_{j}^{+}X(v_{X})$—, and let (see Fig. 2)
$\mathcal{Z}_{j}^{+}=_{\mathsf{def}}\partial(\mathcal{X}_{j}^{+})\setminus(\Psi\big{(}\partial_{j}^{+}X(v_{X})\big{)}\cup\mathcal{Y}_{j}^{+}).$
We claim that, for an odd $j$, $\int_{\mathcal{Z}_{j}^{+}}(d\beta_{Y})^{\frac{2n+1-j}{2}}=0$ since $v_{Y}\in\ker(d\beta_{Y})$ and $\mathcal{Z}_{j}^{+}$ of dimension $2n+1-j$ consists of portions of $v_{Y}$-trajectories.
(a) For an odd $j\in[1,2n-1]$, we get
$\kappa_{j}^{+}(\beta_{X})=\int_{\partial^{+}_{j}X(v_{X})}\big{(}d\beta_{X}\big{)}^{\frac{2n+1-j}{2}}=$
$=\int_{\partial^{+}_{j}X(v_{X})}\Psi^{\ast}\big{(}d\beta_{Y}\big{)}^{\frac{2n+1-j}{2}}=\int_{\Psi\big{(}\partial^{+}_{j}(X)(v_{X})\big{)}}\big{(}d\beta_{Y}\big{)}^{\frac{2n+1-j}{2}}.$
By Stokes’ theorem, applied to $\mathcal{X}_{j}^{+}$, and using that $d\big{(}(d\beta_{Y})^{\frac{2n+1-j}{2}}\big{)}=0$ and that $\int_{\mathcal{Z}_{j}^{+}}(d\beta_{Y})^{\frac{2n+1-j}{2}}=0$, we get
$\kappa_{j}^{+}(\beta_{X})=\int_{\Psi\big{(}\partial^{+}_{j}(X)(v_{X})\big{)}}(d\beta_{Y})^{\frac{2n+1-j}{2}}=-\int_{\mathcal{Y}_{j}^{+}}(d\beta_{Y})^{\frac{2n+1-j}{2}}.$
This validates the claim of Theorem 6.3 for odd $j$.
(b) For an even $j\in[2,2n]$, by Stokes’ theorem, we get
$\kappa_{j}(\beta_{X})=\int_{\partial_{j}X(v_{X})}\beta_{X}\wedge(d\beta_{X})^{\frac{2n-j}{2}}=\int_{\partial_{j-1}^{+}X(v_{X})}(d\beta_{X})^{\frac{2n+2-j}{2}}$
$=\int_{\partial_{j-1}^{+}X(v_{X})}\Psi^{\ast}\big{(}(d\beta_{Y})^{\frac{2n+2-j}{2}}\big{)}=\int_{\Psi\big{(}\partial^{+}_{j-1}(X)(v_{X})\big{)}}(d\beta_{Y})^{\frac{2n+2-j}{2}}.$
Now we are facing the situation, described in the case (a). Therefore,
$\kappa_{j}(\beta_{X})=-\int_{\mathcal{Y}_{j-1}^{+}}(d\beta_{Y})^{\frac{2n+2-j}{2}},$
and thus is determined by $\mathcal{Y}_{j-1}^{+}$, the $\Pi_{Y}$-projection of the locus $\Psi\big{(}\partial^{+}_{j-1}(X)(v_{X})\big{)}$. ∎
Example 6.3.
Let $\dim X=\dim Y=5$. We adopt the notations and hypotheses of Theorem 6.3. Then $\kappa_{1}(\beta_{X})=0=\kappa_{3}(\beta_{X})$ and
[…]
In contrast, we do not know whether $\kappa^{+}_{2}(\beta_{X})$ and $\kappa^{+}_{4}(\beta_{X})$ can be recovered in terms of the $\Pi_{Y}$-generated boundary data residing on $\partial_{1}Y$. $\diamondsuit$
Combining Proposition 6.1 with Theorem 6.3 leads to the following claim.
Corollary 6.2.
Let $X$ and $(Y,\beta_{Y})$ be as in Theorem 6.3. Consider a $t$-family $\{\Psi_{t}:X\hookrightarrow\mathsf{int}(Y)\}_{t\in[0,1]}$ of orientation-preserving contact embeddings such that, for an even $j\in[2,2n]$, the $t$-family of forms $\big{\{}\Psi_{t}^{\ast}\big{(}\beta_{Y}\wedge(d\beta_{Y})^{\frac{2n-j}{2}}\big{)}\big{\}}$ is variationally boundary $(2n+1-j)$-exact in the sense of Definition 6.5.
Then $\int_{\mathcal{Y}_{j-1}^{+}(t)}(d\beta_{Y})^{\frac{2n+2-j}{2}}=-\kappa_{j}(\Psi_{t}^{\ast}\big{(}\beta_{Y}))=-\kappa^{+}_{j-1}(\Psi_{t}^{\ast}\big{(}\beta_{Y}))$ is $t$-independent. $\diamondsuit$
7. Shadows of Legandrian submanifolds of $(X,\beta)$ on the screen $\partial_{1}^{+}X(v_{\beta})$
Let $\dim X=2n+1$. We consider a cooriented contact structure $(X,\xi)$, its contact form $\beta$, and its traversing (boundary generic) Reeb vector field $v_{\beta}$.
We denote by $\mathcal{I}_{x}(d\beta)$, $x\in X$, an isotropic space of $d\beta|_{\xi_{x}}$, that is, a maximal vector subspace of $\ker_{x}(\beta)$ such that $d\beta(w_{1},w_{2})=0$ for any $w_{1},w_{2}\in\mathcal{I}_{x}(d\beta)\subset\ker_{x}(\beta)$. For a contact form $\beta$, $\dim(\mathcal{I}_{x}(d\beta))=n$.
By definition, $L\subset X$ is sub-Legendrian manifold, if $\beta|_{L}\equiv 0$; thus the tangent space $T_{x}L\subset\ker_{x}(\beta)$ is contained in some isotropic space for all $x\in L$.
Remark 7.1.
Certain properties of a contact pair $(X,\beta_{X})$ are inherited by any contact pair $(Y,\beta_{Y})$ that incapsulates $(X,\beta_{X})$ so that $\beta_{Y}|_{X}=\beta_{X}$. For example, if $(X,\beta_{X})$, $\dim X=2n+1$, contains a Legandrian sphere $S^{n}$ or a Legandrian torus $T^{n}$, then so does $(Y,\beta_{Y})$. If $(X,\beta_{X})$ contains a Legandrian sphere $S^{n}$ that bounds an embedded ball $D^{n+1}$, then so does $(Y,\beta_{Y})$. In particular, if a $3$-fold $(X,\beta_{X})$ contains a Legandrian loop $L$ that bounds a $2$-ball $B^{2}\subset X$ (such $\beta_{X}$ is called overtwisted [E]), then any contact $3$-fold $(Y,\beta_{Y})$ that incapsulates $(X,\beta_{X})$ is also overtwisted. $\diamondsuit$
Our next goal is to describe the images $L^{\dagger}$ of sub-Legendrian compact submanifolds $L\subset X$ ($\dim(L)\leq n$) both in $\partial_{1}^{+}X(v_{\beta})$ under the $(-v_{\beta})$-flow and in the trajectory space $\mathcal{T}(v_{\beta})$ under the obvious quotient map $\Gamma:X\to\mathcal{T}(v_{\beta})$. Note that these shadows of $L$ do not change when we apply a $t$-family ($t$ being sufficiently small) of local diffeomorphisms $\phi^{t}$ (as in Definition 4.4), generated by the vector field $v_{\beta}$. Since $\beta$ and $d\beta$ are invariant under $\phi^{t}$, we conclude that $\phi^{t}(L)$ is also sub-Legendrian.
By Lemma 4.3, $\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))$ is an open symplectic manifold, with the symplectic form $\omega^{\dagger}=_{\mathsf{def}}d\beta|_{\partial_{1}^{+}X(v_{\beta})}$. Recall that the form $\omega^{\dagger}$ degenerates along the boundary $\partial_{2}X(v_{\beta})=\partial(\partial_{1}^{+}X(v_{\beta}))$, since there $v_{\beta}\in\ker(d\beta)\cap T(\partial_{1}X)$.
Let $\mathcal{X}_{2}(v_{\beta})$ denote the union of the downward $(-v_{\beta})$-trajectories through the locus $\partial_{2}X(v_{\beta})$, and $L(v_{\beta})$ the union of the downward $(-v_{\beta})$-trajectories through the points of a sub-Legandrian submanifold $L\subset X$. Put $L^{\dagger}=_{\mathsf{def}}L(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})$ (see Fig. 3).
Definition 7.1.
We say that a sub-Legandrian manifold $L\subset X$ is in a weakly general position to the $v_{\beta}$-flow, if $L\setminus(L\cap\mathcal{X}_{2}(v_{\beta}))$ is dense in $L$. $\diamondsuit$
Figure 3. The $(-v_{\beta})$-shadow $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ of a closed Legandrian manifold $L\subset X$, shown as a bold loop in the interior of the bulk $X$. In the figure, $L^{\dagger}$ consists of two arcs. The surface $L(v_{\beta})$ consists of $(-v_{\beta})$-trajectories falling down from $L$. Note the locus $\delta L(v_{\beta})\subset X$ that consists of $(-v_{\beta})$-trajectories falling down from $L^{\dagger}\cap\partial_{2}^{+}X(v_{\beta})$.
Lemma 7.1.
The locus $\mathsf{int}(L^{\dagger})=L^{\dagger}\setminus(L^{\dagger}\cap\mathcal{X}_{2}(v_{\beta}))$ is an open (possibly disconnected) Lagrangian submanifold of the open $(2n)$-manifold
$\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})).$
That is, for any $x\in\mathsf{int}(L^{\dagger})$, the tangent space $T_{x}(\mathsf{int}(L^{\dagger}))\subset\mathcal{I}_{x}\big{(}\omega^{\dagger}\big{)}$, an isotropic subspace of the symplectic form $\omega^{\dagger}=d\beta|_{\mathsf{int}(\partial^{+}_{1}X(v_{\beta}))}$ at $x$.
Proof.
By Lemma 4.3, the $2$-form $\omega^{\dagger}=d\beta|_{\partial^{+}_{1}X(v_{\beta})}$ is symplectic on $\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))$. Since $d\beta$ is $v_{\beta}$-invariant, for any $x\notin\partial_{2}X(v_{\beta})$ and for any isotropic space $\mathcal{I}_{x}\big{(}d\beta\big{)}\subset\ker_{x}(\beta)$, its isomorphic image under the differential of the $(-v_{\beta})$ projection on $\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))$, is an isotropic space for the $2$-form $\omega^{\dagger}$. Therefore, $\mathsf{int}(L^{\dagger})$ is a Lagrangian submanifold of the open manifold $\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$. ∎
Remark 8.1. Since $L$ is embedded in $X$ and $v_{\beta}$ is transversal to $\xi$, $\mathsf{int}(L^{\dagger})$ is immersed in the open $(2n)$-manifold $\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$. $\diamondsuit$
Lemma 7.2.
Let $\gamma$ be a Reeb $v_{\beta}$-trajectory through a point $a_{\star}\in X$, which hits $\partial_{1}^{+}X(v_{\beta})$ transversally at a point $b_{\star}$. Let $L^{\dagger}$ be a given Lagrangian submanifold (with respect to the symplectic form $\omega^{\dagger}=_{\mathsf{def}}d\beta|_{\mathsf{int}(\partial^{+}_{1}X(v_{\beta}))}$) of $\partial_{1}^{+}X(v_{\beta})$, which contains $b_{\star}$. Then, in the vicinity of $a_{\star}$, there exists a unique germ of Legendrian submanilold $L\subset X$ that contains $a_{\star}$ and whose $(-v_{\beta})$-governed projection is the germ of $L^{\dagger}$ at $b_{\star}$.
Proof.
In the vicinity of $a_{\star}$, by the Darboux theorem, we can pick coordinates $(z,x_{1},y_{1},\;\ldots,\;\hfill\break x_{n},y_{n})$ such that: they vanish at $a_{\star}$, $\beta=dz+\sum_{i=1}^{n}x_{i}\,dy_{i}$, $d\beta=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}$, and $v_{\beta}=\partial_{z}$. In these coordinates, the germ at $a_{\star}$ of the Legendrian submanilold $L$ in question is given, in its parametric form, by some smooth functions $z(\vec{t}),\,x_{1}(\vec{t}),\,y_{1}(\vec{t}),\,\ldots\,,x_{n}(\vec{t}),\,y_{n}(\vec{t}),$ where $\vec{t}=(t_{1},\ldots,t_{n})\in\mathbb{R}^{n}$. We assume that the tangent to $L$ vectors
$\tau_{i}=\Big{(}\frac{\partial z}{\partial t_{i}},\frac{\partial x_{1}}{\partial t_{i}},\frac{\partial y_{1}}{\partial t_{i}},\,\ldots,\frac{\partial x_{n}}{\partial t_{i}},\frac{\partial y_{n}}{\partial t_{i}}\Big{)},$
$1\leq i\leq n$, are linearly independent in the vicinity of $a_{\star}$. Put
$\tilde{\tau}_{i}=\Big{(}\frac{\partial x_{1}}{\partial t_{i}},\frac{\partial y_{1}}{\partial t_{i}},\,\ldots,\frac{\partial x_{n}}{\partial t_{i}},\frac{\partial y_{n}}{\partial t_{i}}\Big{)}.$
We first will prove the assertion of the lemma for the case when $\partial_{1}^{+}X(v_{\beta})$ is equal to the hyperplane $H=\{z=const\}$, where the point $\gamma\cap H$ resides in the interval of $\gamma$, bounded by $a_{\star}$ and $b_{\star}$.
The property of $L$ being Legendrian can be expressed by the set of equations:
(7.1) $\{\beta(\tau_{i})=0\}_{1\leq i\leq n}=\Big{\{}-\frac{\partial z}{\partial t_{i}}=\sum_{q=1}^{n}x_{q}\frac{\partial y_{q}}{\partial t_{i}}\Big{\}}_{1\leq i\leq n},$
If these equations are solvable, then the system
(7.2) $\{d\beta(\tau_{i},\tau_{j})=d\beta(\tilde{\tau}_{i},\tilde{\tau}_{j})=0\}_{1\leq i<j\leq n},$
must be solvable as well.
We can solve the equations in (7.1) for $z(\vec{t})$, provided that the RHS of equation (7.1) is a gradient vector field whose components are $\Big{\{}\sum_{q=1}^{n}x_{q}\frac{\partial y_{q}}{\partial t_{i}}\Big{\}}_{1\leq i\leq n}$, or that
(7.3) $\Big{\{}\frac{\partial\tilde{\tau}_{i}}{\partial t_{j}}=\frac{\partial\tilde{\tau}_{j}}{\partial t_{i}}\Big{\}}_{1\leq i<j\leq n}.$
Moreover, the solution $z(\vec{t})$ of (7.1) with the property $z(\vec{0})=a_{\star}$ will be unique.
By a direct computation, the second group of equations (7.2) coincides with the group of equations in (7.3). In turn, the equations in (7.3) claim that $\{d\beta(\tilde{\tau}_{i},\tilde{\tau}_{j})=0\}_{1\leq i\neq j\leq n}$. In other words, they claim that the vectors $\{\tilde{\tau}_{i}\}_{1\leq i\leq n}$ belong to some $n$-space $\mathcal{I}so(d\beta|_{H})\subset H$, so that the $n$-manifold $L_{H}^{\dagger}$, given by the functions $\{x_{1}(\vec{t}),\,y_{1}(\vec{t}),\,\ldots\,,x_{n}(\vec{t}),\,y_{n}(\vec{t})\}$ is Lagrangian in the hypersurfuce $H$. The same argument shows that the germ of any Lagrangian submanifold $L_{H}^{\dagger}\subset H$, passing through the point $\vec{0}\in H$, is the shadow, under the $z$-projection, of the germ of unique Legandrian $L$, containing the origin of $\mathbb{R}^{2n+1}$.
It remains to prove that any germ of a Lagrangian $L_{H}^{\dagger}\subset H$ (with respect to $d\beta|_{H}$) at the point $\gamma\cap H$, with the help of $(-v_{\beta})$-flow, projects onto a germ of a Lagrangian $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ (with respect to $d\beta|_{\partial_{1}^{+}X(v_{\beta})\setminus\partial_{2}X(v_{\beta})}$) at the point $b_{\star}=\gamma\cap\partial_{1}^{+}X(v_{\beta})$. The latter observation follows from the transversality of $\gamma$ to $\partial_{1}^{+}X(v_{\beta})$ at $b_{\star}$ and the $v_{\beta}$-invariance of the form $d\beta$. ∎
For a given Lagrangian submanifold $L^{\dagger}\subset\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))$, consider the (discontinuous) function $\Psi^{\beta}_{L^{\dagger}}:L^{\dagger}\to\mathbb{R}_{+}$, defined by
$\Psi^{\beta}_{L^{\dagger}}(x)=\int_{[x,\;C_{v_{\beta}}(x)]}\beta,$
where $C_{v_{\beta}}$ is the causality map. Note that, for a connected component $L^{\ddagger}$ of $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$, the function $\Psi^{\beta}_{L^{\ddagger}}=\Psi^{\beta}_{L^{\dagger}}\big{|}_{L^{\ddagger}}$ is continuous and smooth.
Put $\mathcal{D}_{L^{\ddagger}}^{\beta}=\sup_{x\in L^{\ddagger}}\big{\{}\Psi^{\beta}_{L^{\ddagger}}(x)\big{\}}.$ One may think of this number as the “Reeb elevation” of the component $L^{\ddagger}$. Evidently, $\mathsf{diam}_{\mathcal{R}}(\beta)\geq\mathcal{D}^{\beta}_{L^{\ddagger}}$.
Let us revisit a given Legandrian manifold $L\subset X$ whose $(-v_{\beta})$-projection is $L^{\dagger}$. For any $x\in\mathsf{int}(L^{\dagger})$, take the first point $y(x)\in\gamma_{x}$ that belongs to $L$ and consider the function
$\psi^{\beta}_{L}(x)=_{\mathsf{def}}\int_{[x,y(x)]}\beta,$
where $[x,y(x)]\subset\gamma_{x}$ is the segment of the trajectory $\gamma_{x}$, bounded by $x$ and $y(x)$.
Evidently, $\psi^{\beta}_{L}(x)\leq\Psi^{\beta}_{L^{\ddagger}}(x)$ for any $x\in L^{\ddagger}\subset L^{\dagger}$.
Corollary 7.1.
Let $\beta$ be a contact form whose Reeb vector field $v_{\beta}$ is traversing and boundary generic. Let $L\subset X$ be a Legandrian submanifold in a weakly general position to “the waterfall” $\mathcal{X}_{2}(v_{\beta})$, the set of $(-v_{\beta})$-trajectories through the locus $\partial_{2}X(v_{\beta})$. Let $L^{\dagger}$ be its $(-v_{\beta})$-guided projection on $\partial_{1}^{+}X(v_{\beta})$.
Consider the connected components $\{L^{\ddagger}_{\kappa}\}_{\kappa}$ of the Lagrangian manifold $\mathsf{int}(L^{\dagger})$, immersed in $\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})).$ In each component $L^{\ddagger}_{\kappa}$, we pick a base point $b_{\kappa}$, distinct from the points of self-intersection of $L^{\dagger}$.
Then the knowledge of the shadow $L^{\dagger}$, together with the set the base points $\{b_{\kappa}\}_{\kappa}$ and the set of values $\{\psi^{\beta}_{L}(b_{\kappa})\}_{\kappa}$, makes it possible to reconstruct the Legandrian embedding $L\subset X$.
Proof.
The reconstruction of $L$ resembles the process of analytic continuation.
Let $\pi_{v_{\beta}}:L\to L^{\dagger}$ be the $(-v_{\beta})$-guided projection. Let $L^{\dagger}_{\times}\subset L^{\dagger}$ denotes the subset consisting of points that have multiple $\pi_{v_{\beta}}$-preimages in $L$, and let $\mathcal{L}^{\dagger}_{2}=L^{\dagger}\cap\mathcal{X}_{2}(v_{\beta})$. By the hypothesis, the set $L^{\dagger}\setminus(L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2})$ is open and dense in $L^{\dagger}$.
Knowing the point $b_{\kappa}\in L^{\ddagger}_{\kappa}$ makes it possible, with the help of the value $\psi^{\beta}_{L}(b_{\kappa})$, to locate a unique point $a_{\kappa}\in L\cap\gamma_{b_{\kappa}}$. There is a unique germ $L_{\kappa}$ of $L$ in the vicinity of $a_{\kappa}$ that projects onto the germ of $L^{\ddagger}_{\kappa}$ at $b_{\kappa}$. Let $U^{\dagger}_{\kappa}$ be an open neighborhood of $b_{\kappa}$ in $L^{\dagger}$ over which the reconstruction of $L$ is achieved.
Take a point $x\in L^{\dagger}$ that belongs to the boundary of $U^{\dagger}_{\kappa}$ and such that $x\notin L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2}$. Then by Lemma 7.2, there exists a unique lift $L^{\bullet}$ to $X$ of $L^{\dagger}$ over some open neighborhood $U_{x}\subset L^{\dagger}$ of $x$, such that the point $y=\gamma_{x}\cap L$ belongs to $L^{\bullet}$. Since, over $U_{x}\cap U^{\dagger}_{\kappa}$, both $L^{\bullet}$ and $L$ are the lifts of $L^{\dagger}$ that share the point $y$, again by Lemma 7.2, we conclude that $L^{\bullet}=L$ over $U_{x}$. As a result, the reconstruction of $L$ over $U_{x}\cup U^{\dagger}_{\kappa}$ is available. Therefore, the maximal open set in $L^{\dagger}$, over which the lift of $L^{\dagger}$, subject to the constraint “$a_{\kappa}$ belongs to the lift”, is possible is a connected component $L^{\ddagger}$ of the open set $L^{\dagger}\setminus(L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2})$. Since $L$ is in weakly general position with $\mathcal{X}_{2}(v_{\beta})$, the union of such lifts, taken over all $\kappa$’s, is a dense subset $L\setminus(L\cap\pi_{v_{\beta}}^{-1}(\mathcal{L}^{\dagger}_{2}\cup L^{\dagger}_{\times}))$ of $L$. Therefore, taking the closure of this union, produces the given $L\subset X$. ∎
Remark 8.1. Note that the very existence of Legandrian $L\subset X$ insured that the assembly instructions in the proof of Corollary 7.1 were consistent.
Even when the Reeb vector field $v_{\beta}$ is convex (so the waterfall $\mathcal{X}_{2}(v_{\beta})=\partial_{2}^{-}X(v_{\beta})$) and the $v_{\beta}$-trajectories $\gamma_{x}$ depend continuously on $x\in\partial_{1}^{+}X(v_{\beta})$, we do not know whether any closed Lagrangian submanifold $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ is the $\pi_{v_{\beta}}$-shadow of a Legandrian manifold $L\subset X$. Corollary 7.1 claims only that, if such $L$ exists, then choosing a point $a_{\star}\in L$ makes it unique. It seems that to tackle this question we need to be able to estimate from above the function $\Psi^{\beta}_{L^{\dagger}}(x)$ (or at least the Reeb elevation $\mathsf{elev}_{\mathcal{R}}(L^{\dagger})$ of $L^{\dagger}$) in terms of the form $\beta$ in the vicinity of $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$. $\diamondsuit$
Proposition 7.1.
Let $\beta$ be a contact form on $X$, and $v_{\beta}$ its traversing boundary generic Reeb vector field. Assume that Property $\mathsf{A}$ from Definition 3.1 is valid. Let $\phi:(L,\partial L)\hookrightarrow(X,\partial_{1}X)$ be a Legandrian embedding of a compact manifold $L$.
Then the following boundary-confined data:
  • the causality map $C_{v_{\beta}}:\partial_{1}^{+}X(v_{\beta})\to\partial_{1}^{-}X(v_{\beta})$,
  • the restriction $(f^{\bullet})^{\partial}:\partial_{1}X\to\mathbb{R}$ of the Lyapunov function $f^{\bullet}:X\to\mathbb{R}$ such that $df^{\bullet}(v_{\beta})=1$,
  • the restriction $\beta^{\partial}=\beta||_{\partial_{1}X}$ of the contact form $\beta$,
  • the Lagrangian666with respect to the symplectic form $d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}$ immersion $\mathsf{int}(L^{\dagger})\subset\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$,
  • the set of base points $\{b_{\kappa}\}_{\kappa}$ and the values $\{\psi^{\beta}_{L}(b_{\kappa})\}_{\kappa}$ as in Corollary 7.1,
allow to reconstruct the Legandrian embedding $\phi:L\hookrightarrow X$, up to a diffeomorphism of $X$ that is the identity on $\partial_{1}X$.
Proof.
By Corollary 4.8, the hypotheses in the first three bullets allow for a reconstruction of $(X,v_{\beta},f^{\bullet},\beta)$ up to a diffeomorphism that is the identity on $\partial_{1}X$. With this quadruple in place, the base points $\{b_{\kappa}\in\mathsf{int}(L^{\dagger})\}_{\kappa}$ and the values $\{\psi^{\beta}_{L}(b_{\kappa})\}_{\kappa}$, determine uniquely a point $y(b_{\kappa})\in L$ on each trajectory $\gamma_{b_{\kappa}}$. By Corollary 7.1, the assertion follows. ∎
Lemma 7.3.
Let $X$ be a compact connected smooth $12n+$ 1-dimensional manifold, $\beta$ a contact form, and $v_{\beta}$ its Reeb vector field, which is traversing.
Let $L\subset X$ be a sub-Legandrian $(2k+1)$-dimensional closed submanifold of $X$, where $2k+1\leq n$. We assume that $L$ is in a weakly general position with respect to the waterfall $\mathcal{X}_{2}(v_{\beta})$. We denote by $L(v_{\beta})$ the set of downward trajectories originating in $L$, and by $\delta L(v_{\beta})$ the set of downward trajectories originating in $L$ and hitting the locus $\partial_{2}^{+}X(v_{\beta})$ (see Fig. 3).
Then, for any $k>0$, we get $\int_{L(v_{\beta})}(d\beta)^{k+1}=0,\;\text{ and }\;\quad\int_{L^{\dagger}}\beta\wedge(d\beta)^{k}=0.$
[…]
Proof.
Since $T_{x}L\subset\ker\beta$ we see that $\beta\wedge(d\beta)^{k}|_{L}=0$. Thus, $\int_{L}\beta\wedge(d\beta)^{k}=0$.
Note that, at a generic point $x\in L(v_{\beta})$, the tangent space $T_{x}(L(v_{\beta}))$ is spanned by $v_{\beta}(x)$ and the image of a tangent space $T_{y}L$ under $(-v_{\beta})$-flow that takes $y\in L$ to $x$. Then, $(d\beta)^{k+1}|_{T_{x}(L(v_{\beta}))}=0$, since $v_{\beta}\in\ker(d\beta)$ Therefore, $\int_{L(v_{\beta})}(d\beta)^{k+1}=0$.
Let $\delta L(v_{\beta})$ be the set of downward trajectories originating in $L$ and intersecting $\partial_{2}^{+}X(v_{\beta})$ (see Fig. 3). Again, at a generic point $x\in\delta L(v_{\beta})$, the tangent space $T_{x}(\delta L(v_{\beta}))$ is spanned by $v_{\beta}(x)$ and a $2k$-dimensional subspace $\mathcal{K}_{x}$ of $T_{x}L$. Thus, for $k>0$, $\beta\wedge(d\beta)^{k}|_{T_{x}(\delta L(v_{\beta}))}=0$ since $v_{\beta}\in\ker\beta$ and $d\beta|_{\mathcal{K}_{x}}=0$. Therefore, $\int_{\delta L(v_{\beta})}\beta\wedge(d\beta)^{k}=0$.
The boundary $\partial(L(v_{\beta}))=\delta L(v_{\beta})\cup L^{\dagger}\cup-L$. In fact, $L^{\dagger}$, away from the set $\delta L(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})$ of measure zero, is an immersed submanifold of $\partial_{1}^{+}X(v_{\beta})$.
Recall that $d(\beta\wedge(d\beta)^{k})=(d\beta)^{k+1}$. Therefore, for $k>0$, by the Stokes’ theorem and using that $\int_{L}\beta\wedge(d\beta)^{k}=0$, $\int_{\delta L(v_{\beta})}\beta\wedge(d\beta)^{k}=0$, and $\int_{L(v_{\beta})}(d\beta)^{k+1}=0$, we conclude that $\int_{L^{\dagger}}\beta\wedge(d\beta)^{k}=0$ as well.
In the special case $k=0$, using similar analysis and Stokes’ theorem, we get
[…]
since $\beta(v_{\beta})=1$. ∎
Corollary 7.2.
(A holographic property of sub-Legandrian links in relation to a concave boundary)
Let $\beta$ be a contact form whose Reeb vector field $v_{\beta}$ is traversing and boundary generic on $X$. Let a sub-Legandrian link777a collection of embedded disjoint loops $L\hookrightarrow X$ be in weakly general position with respect to the waterfall $\mathcal{X}_{2}(v_{\beta})$. Let $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ be the image of $L$ under the $(-v_{\beta})$-flow guided projection $X\to\partial_{1}^{+}X(v_{\beta})$.
Then $\int_{L^{\dagger}}\beta<0$, if and only if, there exists a down-flow trajectory $\gamma_{\star}$ that originates at $L$ and hits the concavity locus $\partial_{2}^{+}X(v_{\beta})$.
On the other hand, if $\int_{L^{\dagger}}\beta\geq 0$, then $\int_{L^{\dagger}}\beta=0$, and there is no such trajectory $\gamma_{\star}$. In such a case, the $L^{\dagger}$ is a collection of closed immersed curves.
Proof.
Since $\beta(v_{\beta})=1$, $\int_{\delta L(v_{\beta})}\beta>0$ when $\delta L(v_{\beta})$ has a positive $1$-dimensional measure. This is equivalent to the existence of downward trajectory that connects $L$ to $\partial_{2}^{+}X(v_{\beta})$.
By Lemma 7.3, in the case $k=0$, we get that $\int_{\delta L(v_{\beta})}\beta>0$ if and only if $\int_{L^{\dagger}}\beta<0$. ∎
We conclude Section 8 with a few natural questions.
Question 7.1.
A $C^{\infty}$-small perturbation $\beta^{\prime}$ of a contact form $\beta$ is a contact form. Is a $C^{\infty}$-small perturbation $v^{\prime}$ of a Reeb vector field $v_{\beta}$ a Reeb vector field for some contact form? For some contact form $C^{\infty}$-close to $\beta$? Equivalently, is the map from the space of contact forms to the space of their Reeb fields open? $\diamondsuit$
Question 7.2.
Assuming that a Reeb vector field $v_{\beta}$ is traversing, how to differentiate between tight and overtwisted contact structures (see [E], [BEM] for the definition) in terms of boundary data? $\diamondsuit$
Question 7.3.
Let $\beta$ be a contact form on a compact manifold $X$ with boundary. Assuming the the Reeb vector field $v_{\beta}$ is traversing, does any Lagrangian888with respect to the symplectic form $\omega^{\dagger}=_{\mathsf{def}}d\beta|_{\mathsf{int}(\partial^{+}_{1}X(v_{\beta}))}$ submanifold $L^{\dagger}$ of the open manifold $\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$ is the image, under the $(-v_{\beta})$-directed projection, of a Legandrian manifold $L\subset X\setminus\mathcal{X}_{2}(v_{\beta})$? The same question, assuming that $v_{\beta}$-flow is convex. $\diamondsuit$
Acknowledgments: I am grateful to Yakov Eliashberg and John Etnyre for very enlightening conversations. Many discussions with Gunther Uhlmann helped me to organize this project; it is my pleasant duty to thank Gunther for the much needed encouragement.